### v1

# for Academic Information Seeking

Abstract
Applying large language models (LLMs) for academic API usage shows promise in reducing researchers’ academic information seeking efforts. However, current LLM API-using methods struggle with complex API coupling commonly encountered in academic queries. To address this, we introduce SoAy, a solution-based LLM API-using methodology for academic information seeking. It uses code with a solution as the reasoning method, where a solution is a pre-constructed API calling sequence. The addition of the solution reduces the difficulty for the model to understand the complex relationships between APIs. Code improves the efficiency of reasoning.
To evaluate SoAy, we introduce SoAyBench, an evaluation benchmark accompanied by SoAyEval, built upon a cloned environment of APIs from AMiner. Experimental results demonstrate a 34.58-75.99% performance improvement compared to state-of-the-art LLM API-based baselines. All datasets, codes, tuned models, and deployed online services are publicly accessible at https://github.com/RUCKBReasoning/SoAy.
[Uncaptioned image] A Solution-based LLM API-using Methodology
for Academic Information Seeking
Yuanchun Wang, Jifan Yu§ [^1] Work was done when these authors interned at Zhipu AI., Zijun Yao§, Jing Zhang, Yuyang Xie§, Shangqing Tu§ Yiyang Fu, Youhe Feng, Jinkai Zhang, Jingyao Zhang, Bowen Huang, Yuanyao Li Huihui Yuan, Lei Hou§, Juanzi Li§and Jie Tang§ Renmin University of China      §Tsinghua University      Zhipu AI https://soay.aminer.cn
1 Introduction
Researchers frequently seek academic information such as a paper’s metadata, a scholar’s publication record, and the connections between scholars as part of their routine scholarly activities. The advent of academic information retrieval systems—including DBLP dblp team (2005), CiteSeer Giles et al. (1998), and Google Scholar Noruzi (2005)—and more sophisticated academic mining platforms such as Microsoft Academic Search Sinha et al. (2015) and AMiner Tang et al. (2008), has substantially reduced the complexity of exploring the academic data, which encompasses massive entities like authors, conferences, and papers.
Figure 1: Comparing responses from Chameleon Lu et al. (2023), ToolLLaMA Qin et al. (2023b), and the proposed SoAy, all leveraging AMiner APIs to address the same academic query.
Large Language Models Zhao et al. (2023) (LLMs) unveil the potential to further reduce the effort required for academic information seeking, benefiting from LLMs’ reasoning ability to comprehend intentions expressed in natural language. Enabling LLMs to utilize existing academic retrieval/mining APIs Qin et al. (2023a) tends to be a feasible solution. However, this adaptation is not easy because of two main challenges:
(1) Coupling: Academic information for academic queries is distributed across various storage locations but interconnected internally, resulting in multiple academic APIs responsible for different functions yet interconnected, i.e. coupled. This requires LLM to fully understand the sequential order of calls between these APIs. Existing API-using methods like Chameleon Lu et al. (2023), shown in Figure 1(a), retrieve and execute APIs separately to get results, thus failing to handle such API coupling.
(2) Efficiency: Existing methods such as ToolLLM Qin et al. (2023b) illustrated in Figure 1(b) employ a depth-first search-based decision tree (DFSDT) for step-by-step reasoning, which has the potential to capture the regularities in the sequence of API. However, the excessive steps lead to query times that users may find unacceptable, compared with traditional academic search engines that offer researchers a very fast response.
We propose SoAy, a Solution-based API-using methodology to apply LLMs to for academic information seeking. Our key innovation lies in the new API calling method: solution and code. As shown in Figure 1(c), SoAy first lets the LLM generate a feasible API calling plan, i.e. solution, based on complex user inputs, and then allows the LLM to generate executable API calling code based on the generated solution. The pre-generated solutions can assist the LLM in clarifying the complex coupling relationships between academic APIs when generating subsequent code. Calling APIs through code generation in a single inference process without repeated reasoning can achieve efficiency levels closer to traditional academic search systems. Specifically, the ability for LLM to reliably utilize complex academic APIs in this manner to accomplish academic search tasks benefits from the automatic generation of a training dataset, consisting of (query, solution, code) triplets. The data is created on the AMiner APIs, resulting in 3,960 automated triplets. We retain 4/5 of these triplets, providing valuable resources for fine-tuning existing LLMs or demonstrating LLM in-context learning. This alignment helps bridge the gap between general LLMs and the intricacies of academic API usage.
To evaluate SoAy, we utilize 1/5 of the generated data, comprising 792 (query, solution, code) triplets. These are enhanced through human verification to ensure the quality of the test set. We then introduce SoAyEval, an evaluation methodology that takes into account both solution correctness and code execution accuracy. By employing SoAyEval for automated evaluation alongside human assessment, our experiments demonstrate the effectiveness and efficiency of SoAy in utilizing AMiner APIs compared to state-of-the-art LLM API-using baselines, across both in-context learning and tuning modes. Furthermore, we validate the versatility of SoAy for application in other academic retrieval systems.
The paper contributes: (1) SoAy, a novel LLM adaptation method, addressing challenges like coupling and efficiency in academic information seeking through automatic data generation, teaching LLMs to use academic APIs. (2) SoAyBench, a released benchmark dataset, accompanied by the SoAyEval evaluation method, to assess LLMs’ proficiency in using academic APIs. (3) Demonstrated effectiveness and efficiency of SoAy in the AMiner system, along with its real deployment on an online service as shown in Figure 13.
Figure 2: Illustration of AMiner API Library, API Graph and solution library.
2 Task Formulation
The task for SoAy is to seek appropriate academic information $I$ based on complicated user query $Q$ in an efficient way. This objective tends to be achieved by using academic APIs in an valid and efficient order.
Problem 1
User Intent Parse Based on $Q$, SoAy needs to understand which academic information the user requires, then identify the associations within them and conduct separate queries.
$Q\rightarrow Q_{S},\quad Q_{S}=\{Q_{i}\}_{i=1}^{N}.$ (1)
Traditional academic search systems cannot accomplish this task, thus requiring users to break their requirements $Q$ down into step-by-step queries $Q_{S}$ to type in the search box. SoAy, on the other hand, needs to use LLMs to break the intent down and get the relevant API calling sequence.
Problem 2
Academic Information Aggregation After understanding what information the user truly needs, SoAy’s task is to retrieve relevant information from data sources according to $Q_{S}$ and aggregate and process the results to provide the user with the precise information $I$ they seek.
Traditional information retrieval systems lack this design and instead display all retrieved information to users for them to select what they want in a list-like format. What SoAy needs to do is generate the correct code to call the academic APIs and process the obtained information within a comparable time to those traditional systems.
Figure 3: The overall process of SoAy around one case of an academic question. The main steps of data generation are shown on the left. Two mode of LLM alignment with the automated triplet dataset are on the right.
3 SoAy: Solution-based API-using
We introduce SoAy to adapt general LLMs for academic information seeking, emphasizing the use of existing academic APIs. To achieve this, abundant training data in the form of (academic query, executable API sequence) pairs is essential, facilitating either fine-tuning of existing LLMs or guiding them through in-context learning. Illustrated in the left part of Figure 3, this data generation pipeline comprises four key steps:
(1) Solution Library Construction. Utilizing the provided academic APIs, we construct a graph representing the dependencies among APIs, and traverse this graph to identify valid and simple paths, referred to as solutions. These solutions collectively form a library known as the solution library.
(2) Combination Formulation. To obtain a more precise semantic of each solution in the library, we combine all the possible inputs of the initial API and all the outputs of the final API of the solution to formulate a series of combinations.
(3) Query Generation. We employ ChatGPT to transform each combination into a template question lacking specific academic entity data. These templates are then diversified into various expressions and populated into potential academic queries with data from a specialized academic source, the AMiner database here.
(4) Code Generation. Finally, we use ChatGPT again to write a piece of corresponding code given the combination, which is expected to correspond to the template query that shares the same combination. Lastly, we employ ChatGPT once more to generate the corresponding code for the given combination. This code is anticipated to match the template query that has the same combination.
Notably, in the data preparation part, we indicate ChatGPT as gpt-3.5-turbo-0613. We present the details of the data preparation prompt and the process of setting up SoAyBench in the appendix.
4 Solution Library Construction
Given an academic API Library $L_{A}$, the initial step is to build an API Graph where each node represents an API $f$, and a directed edge from node $f_{n}$ to node $f_{m}$ signifies a coupling relationship from $f_{n}$ to $f_{m}$ with the intersected attributes $a_{n,m}$.
Definition 1
API Coupling defines the frequent dependencies of the inputs and outputs in academic API calling, caused by academic entity correlations’ complexity:
$f_{n}$ $\xrightarrow{a_{n,m}}f_{m},$ (2)
$if\quad a_{n,m}=i_{m}$ $\cap o_{n}\quad\text{and}\quad a\neq\emptyset,$
where $i_{n}$ denotes the input parameters of $f_{m}$ and $o_{n}$ denotes the returned attributes of $f_{n}$.
[…]
Definition 2
A Solution $S$ is defined as an execution sequence of APIs traversed from the API graph:
$S=f_{1}\xrightarrow{a_{1,2}}f_{2}\cdots\xrightarrow{a_{k-1,k}}$ $f_{k}\cdots\xrightarrow{a_{J-1,J}}f_{J},$
$\forall k,f_{k}\in$ $L_{A}.$ (3)
A solution library $L_{S}$ denotes a set of valid and simple solutions, ensuring by the following ways.
Valid Solution. Among all traversed solutions from the API graph, validity is determined by starting with a fuzzy API, as only fuzzy APIs can accept user utterances in natural language as input.
Simple Solution. For a single query, there may be multiple solutions available. The preferred choice is the simplest one, containing the fewest APIs. For instance, when inquiring about the publication date of the representative work of a scholar, the optimal solution is “searchPerson $\rightarrow$ getPersonPubs”, despite obtaining the same information is also possible through “searchPerson $\rightarrow$ getPersonPubs $\rightarrow$ getPublication $\rightarrow$ searchPublication”.
We meticulously craft prompts to instruct ChatGPT in fulfilling the above two characteristics. All prompts for question generation are detailed in Appendix Part III. Prompts. Consequently, we’ve created 16 solutions to establish the solution library $L_{S}$ for AMiner APIs. Figure 2(c) presents a subset of these solutions, including their input parameters, returned attributes, and corresponding question templates.
5 Reasoning Code Generation
In order for LLM to aggregate effective academic information and provide users with precise answers, SoAy let LLM generate code containing API calls for academic information retrieval. To achieve this, a large amount of Query, Corresponding Code data needs to be generated to train LLM.
5.1 Combination Formulation
Section 2 describes the one-to-several relationship between a query and multiple solutions. It’s evident that a similar one-to-several relationship exists between a solution and several query intents. In a solution $S$, the first API is referred to as the head API $f_{1}$, while the last API is termed the tail API $f_{J}$. For example, in Figure 2(c), consider the 3-hop solution “searchPublication $\rightarrow$ getPublication $\rightarrow$ getPersonBasicInfo”. Due to the diversity of the Head API Input $i_{1}$ and the Tail API Output $o_{J}$, determining which information (outputs) is sought under what conditions (inputs) remains uncertain even after the solution is established. An intuitive approach is to combine the head API input, the solution, and the tail API output to refine the intent space into a specific requirement. A solution-based combination is defined as a specific binding of the head API input, the solution, and the tail API output: $C=(i_{1},S,o_{J})$. Given the $i_{1}$ is “title” and the $o_{J}$ is “education_experience”, the $S$ “searchPublication $\rightarrow$ getPublication $\rightarrow$ getPersonBasicInfo” could lead to one plausible interpretation: querying the graduation school of the first author of the provided publication. It’s important to recognize that the information from the academic data source corresponding to $i_{1}$ is both the information needed by q-t to fill and the input information needed by the generated code, $o_{J}$ is both the expected answer of q-t and the golden result of the code.
5.2 Query Generation
Template-Query Generation. For a given combination $C$, we define a template question $Q_{T}$ to represent the intent of $C$ with missing entity information. In the example provided, the corresponding template question $Q_{T}$ could be formulated as “Which school did the first author of XXXX graduate from?”, where “XXXX” refers to a placeholder representing the title of a publication, as illustrated in Table 1.
Semantic Expansion and Entity Filling. A specific intent represented by a template question $Q_{T}$ may be expressed by multiple natural language queries. For instance, “Which school did XXX graduate from?” and “What is XXX’s alma mater?” convey the same intent. To enrich the semantic diversity of an intention, we conduct semantic augmentation by crafting prompts to leverage ChatGPT in generating three new, semantically enriched template questions for each combination.
Following semantic augmentation, we replace the special symbols in the templates (represented by uppercase “X”s) with information about a specific academic entity in an academic data soruce. A placeholder in a question template must be replaced by the attributes of specific academic entities, such as person_id, name, interest, organization, bio, educational_experience, etc., as illustrated in Figure 2. Taking the instance of the alma mater query again, The name of this scholar will be used to fill in the $Q_{T}$ to become a real question “What is scholar name’s alma mater” and the educational experience of this scholar serves as the golden answer to this query.
Table 1: The semantics of symbols in $Q_{T}$.
Symbol Semantics
X interest
XX organization
XXX name
XXXX publication_info
N count or numbers
Y Year
5.3 Code Generation
In the final step, we employ ChatGPT again to generate a piece of code that corresponds to a given combination $C$ with missing input information. This code is intended to get the answer of the template query $Q_{T}$ shared the same combination $C$ with it. Still with information from a specific academic entity, the $i_{1}$ serves as the input of the generated code. The code is expected to be executed to get the same result as $o_{J}$, which can be seen as a golden information of this code. If the result matches the golden information, this code is valid. Any invalid code, not producing matching results, will be excluded from the final dataset.
6 LLMs Alignment
In order to ultimately leverage LLM to accomplish the task of academic information seeking, SoAy needs to align LLM to be capable of calling academic APIs correctly and efficiently through generating solutions and code. This can be achieved by constructing training data and aligning the LLMs according to it, whether through a tuning or tuning-free method.
6.1 SoAyBench–{(Query, Solution, Code)}
Table 2: Question statistics in SoAyBench set.
Question Type One-hop Two-hop Three-hop Total
Scholar 540 1,800 540 2,980
Publication 180 180 720 1,080
Total 720 1,980 1,230 3,960
After going through the aforementioned four stages, a substantial amount of high-quality (Query, Solution, Code) data has been automatically collected. We organize this data into a comprehensive benchmark known as SoAyBench. Currently, SoAyBench consists of 3,960 triplets, out of which one-fifth has undergone expert inspection and adjustment to form a test set. This test set is used to evaluate the ability of the LLMs to search academic information using the AMiner API. The remaining portion of the data is utilized to support the alignment of the LLMs. Table 2 shows the question statistics of SoAyBench.
Table 3: Overall Performance comparison of the proposed SoAy and the baseline methods ToolLLaMA and GTP-DFSDT, as tested on SoAyBench. The terms 3.5, 3.5-16K, and 4 are abbreviations for different versions of GPT. 7B and 13B denotes the parameter scale of the LLM. SoAyBench is derived from 44 question templates, we calculate each metric within a group of instantiated questions of these templates, and then report the mean and variance across groups with different instantiated entities.
[…]
DS(%) WS(%) WP(%) EE(%)
ToolLLaMA 7B one-hop 12.50±8.00 24.31±13.26 1.39±0.00 54.17±16.01 7.64±5.20 20.14 16.72
two-hop 10.10±4.10 47.22±12.28 0.76±2.27 38.13±9.62 3.79±2.92 13.89
three-hop 11.51±6.53 38.10±14.27 1.19±3.57 43.25±13.07 5.95±4.59 17.46
GPT-DFSDT 3.5 one-hop 55.56±21.06 15.28±7.80 4.86±0.00 21.53±10.67 2.78±0.00 58.33 43.22
two-hop 29.55±11.47 34.34±9.23 4.29±3.64 25.76±8.65 6.06±4.11 35.61
three-hop 38.10±15.09 28.57±11.35 3.17±2.50 25.00±8.87 5.16±6.19 43.25
3.5-16k one-hop 25.69±10.91 9.72±5.00 2.78±0.00 22.92±9.47 38.89±15.60 64.58 43.67
two-hop 16.92±7.76 15.91±6.05 3.28±1.31 46.97±7.13 16.92±4.99 33.84
three-hop 18.65±7.37 15.48±5.63 2.78±0.00 38.49±10.43 24.60±8.53 43.25
4 one-hop 27.78±9.60 2.08±0.00 4.17±5.00 28.47±6.82 37.50±10.91 65.28 58.16
two-hop 26.26±8.89 9.60±4.88 17.93±5.40 15.15±5.39 31.06±9.12 57.32
three-hop 22.22±8.65 7.54±4.46 17.06±6.96 19.05±6.45 34.13±9.87 56.35
SoAyGPT 3.5 one-hop 27.78±8.70 15.97±7.73 3.47±0.00 13.19±7.80 39.58±9.12 67.36 67.30
two-hop 33.84±4.94 9.60±4.75 6.06±2.81 13.13±7.12 37.37±5.06 71.21
three-hop 22.22±6.43 12.70±5.91 9.52±4.42 13.10±6.72 42.46±6.00 64.68
3.5-16k one-hop 28.47±11.67 15.28±6.12 1.39±0.00 17.36±7.78 37.50±9.07 65.97 66.76
two-hop 35.86±6.01 7.32±3.41 5.30±2.18 15.91±7.16 35.61±4.65 71.46
three-hop 23.02±7.16 10.32±4.99 8.33±3.42 17.46±7.37 40.87±6.26 63.89
4 one-hop 0.00±0.00 0.00±0.00 1.39±0.00 2.78±0.00 95.83±5.70 95.83 86.57
two-hop 15.91±4.71 1.26±0.00 9.34±1.07 2.02±1.69 71.46±3.74 87.37
three-hop 6.75±0.00 0.40±0.00 14.68±1.68 1.98±0.00 76.19±3.25 82.94
SoAyLLaMA Chat-7B one-hop 0.00±0.00 0.00±0.00 0.00±0.00 0.69±0.00 99.31±2.94 99.31 85.76
two-hop 0.00±0.00 0.00±0.00 20.20±3.84 2.53±1.97 77.27±2.70 77.27
three-hop 0.00±0.00 0.00±0.00 9.92±3.56 3.17±2.50 86.90±2.72 86.90
Code-7B one-hop 0.69±0.00 0.00±0.00 0.69±0.00 5.56±4.37 93.06±7.50 93.75 88.95
two-hop 0.25±0.00 3.28±0.00 7.07±2.75 4.80±3.69 84.60±5.18 84.85
three-hop 0.40±0.00 0.00±0.00 4.76±2.14 5.16±4.57 89.68±6.54 90.08
Code-13B one-hop 0.00±0.00 0.00±0.00 1.39±0.00 0.00±0.00 98.61±4.03 98.61 92.74
two-hop 0.00±0.00 2.27±0.00 14.14±2.14 0.51±0.00 83.08±3.32 83.08
three-hop 0.00±0.00 0.00±0.00 2.38±2.86 0.40±0.00 97.22±4.28 97.22
6.2 Model Alignment
We provide alignment strategy for both super large closed-source LLMs and tunable open-source LLMs. (1) SoAyGPT is our proposed solution for models like GPT Eloundou et al. (2023), which cannot have their parameters modified but possess strong instruction following capabilities. We select a diverse set of triplets from SoAyBench data that represent different usages of solutions as examples to construct prompts. Given a query, SoAyGPT generates the solution and code in sequence through in-context learning, followed by executing the generated code to obtain the result. We utilize different versions of GPT models as backbones for this purpose. (2) SoAyLLaMA, on the other hand, is our alignment strategy for the LLaMA Touvron et al. (2023) model, which is an open-source model that we can fine-tune within an acceptable training cost. By using the query of the triplets from SoAyBench as input and integrating the solution and code as output, SoAyLLaMA is trained to generate the solution and code in a single step based on the given query in a zero-shot setting.
7 Experiments
In this section, we conduct experiments to validate the effectiveness, efficiency, and practicality of SoAy. We focus on the following Research Questions: (RQ1) In which part of the academic information seeking do LLMs struggle the most? (RQ2) In terms of effectiveness on academic information seeking, does SoAy outperform existing methods? (RQ3) In terms of efficiency, does SoAy finish the task within an execution time that users can accept? (RQ4) What kind of base model is more suitable for constructing SoAy?
We conduct experiments on SoAyBench to compare SoAy with other baseline methods, focusing on these research questions. Additionally, we introduced a detailed evaluation method called SoAyEval, which could also be used in the evaluation of other tool-using scenarios. We also perform online evaluations specifically targeting real-world academic information seeking use cases. There’s an independent chapter in the appendix showing more implementation details and interesting findings.
7.1 Experimental Settings
Baselines. Many existing tool-utilizing works fail to function effectively due to API inter-dependency and intricate execution logic. We have chosen GPT-DFSDT, employing the same three versions of the GPT as backbone as SoAyGPT and ToolLLaMA-7b Qin et al. (2023b) as baselines in API-using experiment on SoAyBench, as the former represents the state-of-the-art model on a general domain tool-utilizing benchmark, ToolBench Qin et al. (2023b), and the latter is an open-source model that has been fine-tuned on numerous tool-utilizing data samples, achieving results comparable to GPT-DFSDT. For every LLM, we adopt a temperature setting of 0 to ensure reproducibility.
It is worth noting that the responses from the baselines, ToolLLaMa, and GPT-DFSDT, are in natural language format, whereas the answers in SoAyBench are the responses returned by SoAPI. In order to ensure a fair comparison, we have designed a set of prompts to employ ChatGPT to extract precise answers from responses in a natural language format.
SoAyEval. Given a question that yields an API solution along with an invoked answer, we have outlined five types of evaluation metrics. These metrics take into account the accuracy of the solution, program, and answer together: (EM): Both the retrieved solution and answer Exactly Match the ground truth. (DS): The answer is correct, but a Different Solution is retrieved compared to the ground truth. (WS): The answer is wrong due to a Wrong Solution. (WP): The solution is correct but the answer is wrong, due to a Wrong Program generated for the solution, which can be executed but yields the wrong answer. (EE): Execution Error, which may by caused by the generation of a non-executable program or network errors during the APIs request. Based on the aforementioned metrics, we define the accuracy metric as follows:
[…]
which represents the proportion of correct answers produced by the method under test. Finally, we assign different weights $w_{1}$, $w_{2}$, and $w_{3}$ to the ACC of one-hop, two-hop, and three-hop of the tested method to compute the overall Score:
$\text{Score}=\frac{w_{1}\cdot\text{ACC}_{1}+w_{2}*\cdot\text{ACC}_{2}+w_{3}\cdot\text{ACC}_{3}}{w_{1}+w_{2}+w_{3}}.$
We name this evaluation method SoAyEval. SoAyEval meticulously analyzes the entire process of large models using a set of APIs to answer corresponding questions. It is a universal evaluation method that could be adopted by other API-using scenarios. Notably, we assign $w_{1}=1,w_{2}=2,w_{3}=3$ in our experiment to highlight the multi-hop question-answering ability of the methods under test.
7.2 Overall Performance (RQ1, 2)
Table 3 presents the experimental results of SoAyGPT, SoAyLLaMA, and the other two representative baseline methods on SoAyBench. SoAyLLaMA achieves the highest overall score, with Code-13B Model of 91.24%, followed by the other SoAy method, SoAyGPT, which gets 86.57% with GPT-4 as the backbone. These results prove the effectiveness of SoAy. With the help of the prepared solution library, SoAy describes the complex inter-dependency and diverse execution logic to the LLMs. This significantly reduces the pressure of LLMs when determining the optimal execution path from the coupling APIs.
Despite undergoing specialized tool usage training, ToolLLaMA does not score highly. This is primarily because ToolLLaMA’s training objective is to master the use of general APIs, where the dependencies involved are significantly less complex than those in complicated academic APIs. Furthermore, many questions in SoAyBench require a solution that incorporates branch or loop logic. Facing with the need to execute these complicated execution logics, GPT-DFSDT’s decision-tree-based inference encounters numerous repetitive sub-nodes, making the solution exceptionally challenging.
7.3 Error Analysis (RQ1)
The detailed analysis of model performances from SoAyEval reveals distinct characteristics and potential improvement areas for each model across different question types. The SoAyLLaMA and SoAyGPT models, particularly their latest versions, demonstrate superior capabilities in managing complex question types with minimal errors. ToolLLaMA and early versions of GPT-DFSDT show greater susceptibility to execution and solution pathway errors, indicating potential areas for algorithmic or training enhancements. Overall, advancements in model versions (notably from GPT-DFSDT 3.5 to 4 and SoAyGPT 3.5 to 4) show tangible improvements in reducing specific error types and enhancing overall accuracy, highlighting the benefits of continuous model development and refinement.
7.4 Efficiency (RQ3)
We evaluate the efficiency of the proposed method by reporting the average response time of various methods evaluated on SoAyBench, as shown in Table 4. We apply both SoAyLLaMA and ToolLLaMA in a network request service format to keep justice with GPT-based methods which should only be accessed through APIs calls. Both ToolLLaMa and GPT-DFSDT rely on multiple executions to complete complex API operations, thus consuming a considerable amount of time. Implementation through SoAy can effectively reduce the inference time of tool calls. Even with GPT-4, the average execution time of SoAy is nearly one-third of the execution time of GPT-DFSDT with GPT-4. Moreover, with a smaller parameter scale, the lightweight SoAyLLaMA models show even less execution time, which we consider already meet the requirement for being applied to an academic searching service.
Table 4: Average response time of different methods (second).
Method 7B 13B 3.5 3.5-16k 4
ToolLLaMA 45.10 / / / /
GPT-DFSDT / / 39.12 53.73 70.92
SoAyGPT / / 6.15 6.40 26.05
SoAyLLaMA-Code 1.12 1.35 / / /
7.5 Backbones Comparison (RQ4)
It’s worth considering which backbone model to choose to construct an academic searching system whether through SoAyGPT or SoAyLLaMA. Based on the results of SoAyBench, if we focus on the effectiveness of the results, the choice of SoAyGPT definitely falls on the latest version (GPT-4 when the paper is drafted). As for the choice of SoAyLLaMA, whose implementation is described in Appendix B, we can find a trend where the scores increase as the number of parameters grows. Simultaneously, given that Soay implements the planning and formatting processes defined in the problem section through code generation, CodeLlama, which excels in code generation, outperforms ChatLlama. It should be mentioned that the effectiveness shown on the results of SoAyBench should not be the sole criterion for judgment when considering practical applications.
8 Related Works
APIs, short for Application Programming Interfaces, are the software interfaces that provide services to other software components [^1] https://en.wikipedia.org/wiki/API#cite_note-1. It facilitates the exchange of data between various software applications, thereby enabling interoperability between them Ofoeda et al. (2019). For the application of LLMs using the API-using methodology, the current solutions and technical approaches can be divided into two major groups: A straightforward idea is to broaden the tool/API-using coverage for LLMs. Such efforts not only include basic web services, such as search engines Nakano et al. (2021); Liu et al. (2023); Zhang et al. (2023) and calculators Schick et al. (2023), but also include the intelligent multi-modal functions, such as vision understanding Wu et al. (2023) and image generation Lu et al. (2023).
Others are devoted to the tool’s dependencies, programming logic and planning ability King et al. (2023); Shen et al. (2023); Paranjape et al. (2023). Except for the efforts that regard LLMs as database managers Qin et al. (2023a), there are also plenty of methods that modeling the tool’s graphical relations Zhang (2023), building intermediate instructions Burnett and McIntyre (1995) and accomplish complex tasks Song et al. (2023).
Besides, applying LLMs to directly access the data from the Database through techniques like text-to-SQL Qin et al. (2022) proves ineffective. The limitation arises from the unavailability of access to the original data sources of related systems.
9 Conclusion
In this work, we propose an API-Using methodology, SoAy, to apply LLMs to academic information searching. The core technique of SoAy is the high-quality automatic built dataset, SoAyBench, providing 3960 academic triplets data. We align LLMs to generate codes on pre-constructed guidance, which could face the coupling and execution logic diverse challenges. An evaluation strategy, SoAyEval, is proposed to evaluate the AMiner API-using capability. Experiments show SoAy’s effectiveness, efficiency, and usability. In the future, we will continue to enlarge the SoAyBench and apply SoAy to more scenarios and domains.
10 Limitations and Future Directions
In this paper, we propose an approach that combines automated data generation with LLM alignment for academic query API-based tasks and demonstrates its effectiveness. The released SoAyBench and the proposed SoAyEval could support for evaluating related methods in subsequent studies. However, there are still some worthwhile explorations that we have not conducted:
  • SoAyBench is a large-scale dataset constructed through automation, consisting of 3960 triplet data instances. Although we incorporate quality checks in the solution library construction, query generation, and code generation processes, and have human experts inspect and refine the test set, we must acknowledge that we cannot guarantee the provided tests in SoAyBench align perfectly with people’s expectations for the API system. Some prior worksQin et al. (2023b) validate the quality of their benchmarks by comparing the results of tested methods on their own benchmark with manually evaluated results. This is an aspect we plan to further explore and validate in the future, which may lead to more human-aligned automated constructed datasets.
  • The API-using approach we propose addresses the task of guiding large models to utilize complex academic search APIs. However, we have not yet validated whether our method has advantages in simpler scenarios. We will continue to expand SoAyBench and conduct practicality and effectiveness validations in various scenarios.
  • In practical applications, we have encountered an inconvenient point of updating the deployed system: if we want to incorporate new APIs on top of existing ones, especially when adding a large number of APIs, the context-guided SoAyGPT needs to adjust the structure of the prompt within a limited input space. This poses a challenge that requires a trade-off between the breadth and granularity of examples. Although our proposed training-based SoAyLLaMA theoretically allows for the unlimited expansion of APIs by adding sufficient training data, this is an aspect we have not explored yet. In the future, we plan to explore methods for continuous learning with SoAyLLaMA to enable rapid inclusion of new API data without compromising the model’s mastery of existing APIs.
In conclusion, the quality of data generation, scalability, and the scenarios presented in this paper opens up numerous possibilities for further exploration. In the future, we will continue to fulfill the current work and explore further in these mentioned directions.
11 Acknowledgements
This project would like to thank the data annotators at Zhipu AI for their help and support. All GPU compute and API costs are provided by Zhipu AI.
12 Ethical Considerations
12.1 New datasets
Using SoAy, our new dataset is built on the AMiner dataset. The construction of this new dataset relies on the efforts of annotators. We value their contributions and provide them with reasonable compensation based on the quantity of annotations and the duration of their work. For the characteristics of the dataset, it focuses on the academic field and is suitable for academic researchers. Furthermore, as talked before, to prevent additional risks, we establish an evaluation benchmark, which is further enhanced by additional human verification to ensure the quality of this new dataset and its application effectiveness.
12.2 NLP application
The intended use of the SoAy trained LLMs is to facilitate researchers in better accessing academic literature and other related resources. It allows for querying in natural language, reducing the technical barriers for scholars and others with academic information retrieval needs. However, just like the issues inherent to LLMs, sometimes it may provide incorrect information, potentially leading to misguidance for users, which may be a serious disruption for researchers since academic information needs to be accurate.
Appendix
In the appendix, we’ll present in three folds the details of SoAy and SoAyBench, along with more experimental details and conclusions. Part I focus on the implementation of SoAy, including the details of SoAyGPT in  A, SoAyLLaMA in  B and SoAyBench in  C. Part II shows more about the experiments, including the online evaluation in  4, the implementation of SoAy on other academic information seeking systems in  F, and more interesting findings from the results on SoAyBench in  E. Part III lists all the prompts used in SoAy and SoAyGPT.
Part I. Implementation Details
The content of this section primarily introduces some implementation details of the SoAy method, including the inference architecture of SoAyGPT, the training details of SoAyLLaMA, as well as the data details and API environment setup details of SoAyBench. We implement SoAyGPT on three distinct versions of GPT, specifically 3.5: gpt-3.5-turbo-0613; 3.5-16k: gpt-3.5-turbo-16k-0613, and 4: gpt-4-0613. We conduct the fine-tuning on 3 different variants of the LLaMA model, respectively Chat-7B: Llama-2-7b-chat-hf, Code-7B: CodeLlama-7b-Instruct-hf, and Code-13B: CodeLlama-13b-Instruct-hf.
Appendix A SoAyGPT
SoAyGPT is an alignment method designed for a series of large-scale models that are closed-source yet possess high command-following capabilities. Specifically, SoAyGPT employs in-context learning to construct three LLM agents Wang et al. (2024) (Solution Retrieval Agent, Code Generation Agent, and Answer Generation Agent) and an environment operable by these agents to complete the alignment process. The agents and environments are designed:
Solution Retrieval Agent. Given the user input $Q$, the solution retrieval agent obtains the appropriate solution from the solution library $S$. It’s important to note that the retrieval process here diverges from traditional information retrieval functions, which extract the exact required element from a set. Essentially, this retrieval module presents all possible solutions as the prompt to the LLMs, enabling them to generate a suitable solution. Consequently, it’s possible that the produced solution differs from any solution in $L_{S}$.
Code Generation Agent. The code generation agent must generate the correct program under the guidance of the retrieved solution $S$. Since the solution might not be included in $L_{S}$, if the retrieved solution $S$ is in $L_{S}$, we simply use $P_{p}$ specifically defined for $S$ to generate the required program. Otherwise, SoAy has to use $P_{g}$ that is defined for all the possible solutions to guide the LLMs in generating a program, as no specific solution could be identified. SoAy employs Python as the programming language for each program, allowing it to incorporate not only sequential but also branching and looping logic.
Code Execution Environment. SoAy employs a virtual Python environment as the executor to obtain the results, serving as the knowledge $K$. Here, $K$ is expected to be the exact answer to $Q$ if $Q$ asks about some precise information, like the number of citations for a scholar. Otherwise, $K$ provides helpful information if $Q$ has an ambiguous intent.
Answer Generation Agent. Finally, SoAy generates a natural language answer $A$ based on $Q$ and $K$, guided by $P_{a}$.
Appendix B SoAyLLaMA
SoAyLLaMA is designed for supervised fine-tuning of open-source models when training resources are available. Through the training stage, SoAyLLaMA integrates an API description and queries as inputs and combines solutions and programs as responses. During the inference stage, SoAyLLaMA concurrently completes the Planning and Formatting processes to derive solutions and code. Similar to SoAyGPT, it utilizes the environment to execute and obtain answers based on $K$. We conduct the fine-tuning on 3 different variants of LLaMA Model, respectively Chat-7B: Llama-2-7b-chat-hf [^2] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, Code-7B: CodeLlama-7b-Instruct-hf [^3] https://www.modelscope.cn/models/AI-ModelScope/CodeLlama-7b-hf/summary, Code-13B: CodeLlama-13b-Instruct-hf [^4] https://huggingface.co/meta-llama/CodeLlama-13b-Instruct-hf.
Dataset. As mentioned in Section 3, four-fifth of the instances from SoAyBench are utilized here to train the SoAyLLaMA. We simply take the academic query as an input and the relevant solution.
Training Details. We train 13B models on 8 * NVIDIA A100 GPU (80GB) for 13 hours to accomplish a convergent loss, with a batch size of 8. 7B models are trained on 4 * NVIDIA A100 GPU (80GB) for 7 hour with a batch size of 16. All the training details are saved in the training log in our Github Repo [^5] https://anonymous.4open.science/r/SoAy-C025/readme.md.
Appendix C SoAyBench
To assess AMiner API-using capabilities, the foundational AMiner APIs for LLMs to invoke and a test set should be provided. However, given the dynamic and specialization nature of academic data, with academic entities-related information rapidly changing, maintaining a test set with static answers proves challenging. To address this challenge, we clone AMiner’s APIs at a specific point in time to create a static service, from which we create a corresponding static test set including 792 meaningful academic questions (one-fifth of the whole SoAyBench) and ground truth conducted from the cloned APIs. We extract a subset of the comprehensive AMiner database, which includes over 66,000 entities such as persons and publications, at a specific point in time, September 23, 2023. This subset serves as our experimental evaluation environment, and we clone an API Service based on it. Consequently, the results returned by these APIs, given the same parameters, will remain consistent over time. This dataset and the API service provide a foundation for our subsequent construction of the test set.
Part II. Experiments Details
In this section, we will present more details related to the experiments. To verify the efficiency and effectiveness of SoAy in real applications, it is insufficient to rely solely on its performance on SoAyBench, so we designed an online evaluation to test SoAy’s performance in real-use scenarios. Beyond the research questions in Section 7, more interesting insightful conclusions can be found in the results on SoAyBench. Furthermore, to validate the universality of the SoAy method, rather than it only working for the AMiner system, we implemented SoAyGPT in two other academic information seeking scenarios and demonstrated its results in a case study manner.
Appendix D Online Evaluation
Figure 4: Distribution of questions used for online human evaluation. Questions are categorized by 1, 2, and 3-hop types, further divided into exact and ambiguous categories. The ambiguous category is specific to this online evaluation setting.
We examine the Spearman correlation coefficients among the voting results of the 15 human annotators, as shown in Figure 5. The released SoAyBench contains the templated questions derived from the AMiner SoAPI, thus testing a method’s ability to understand and utilize the SoAPIs to a certain extent. However, in real-world deployment scenarios, the goal of a good method should be to fulfill user needs more comprehensively and offer a superior user experience. To achieve this, we implement the proposed SoAy as an online application, gather 52 real user demands from the logs, and invite 10 annotators to conduct human evaluation.
Figure 5: Spearman correlation among 15 annotators
Settings. We deploy SoAy with GPT-3.5-16k as its backbone online, due to the superior speed of the 3.5 version, enhancing user experience. Over time, we gather 52 authentic questions from the application logs. The distribution of these questions is illustrated in Figure 4, categorized not only by their 1, 2, and 3-hop nature but also classified as either exact or ambiguous. Ambiguous questions, such as requests for “comparisons between two scholars”, lack standardized answers and thus yield diverse responses. We find that the majority of these 52 questions could be resolved within two hops.
To evaluate the answers, we invite five academic experts to provide responses based on their expertise, supplemented by the AMiner system and general search engines. The online system’s answers, responses directly from ChatGPT, and the aforementioned human-provided answers constitute the comparison methods.
For a balanced and fair assessment of the experimental results, we develop an annotation platform and create a detailed annotation manual. This platform allows annotators to shuffle and anonymize the answers from the three methods before each voting session. A team of 10 annotators is recruited to assess the answers provided by each method for every question across a two-day period. Their focus is on evaluating the accuracy and readability of the responses, culminating in a vote for the most effective response among the three. We also open-source the platform and its associated manual.
Evaluation Results.
Figure 6: Voting results for authentic online questions.
After identifying the three annotators with the lowest correlation to others, we exclude their input. For the remaining annotators, a majority voting strategy is employed to consolidate the votes for the three methods. The summarized results are presented in Figure 6. These results show a preference for the SoAy method’s responses in real-world scenarios across various question types, outperforming both direct ChatGPT and human expert responses.
Notably, SoAy’s responses are particularly favored for questions with exact answers. However, SoAy’s advantage is less pronounced for ambiguous questions when compared to baseline responses. This observation aligns with SoAy’s design and the assessment objectives of SoAyBench, both of which prioritize precise answering through the integration of SoAPI Library resources. Nevertheless, there’s a noticeable gap between the nature of real-world inquiries and the templated questions in SoAyBench, posing a challenge for SoAy in addressing certain scenarios. This finding points towards the potential for integrating SoAPI calls with model-intrinsic functions for a more robust response system.
Appendix E Insights on SoAyBench
E.1 Intriguing phenomena
Table 3 reveals several other intriguing phenomena: First, while SoAy excels with GPT-4, its advantages are less marked on GPT-3.5-turbo and GPT-3.5-16k (hereafter collectively referred to as GPT-3.5 backbones), trailing behind the GPT-DFSDT method with the same GPT-3.5 backbones. Notably, the DS and WS metrics for SoAy with GPT-3.5 backbones are quite high. This can be attributed to GPT-4’s superior instruction-following proficiency, which amplifies the impact of SoAy’s complex prompts on the model adept at instruction following. Second, SoAy methods exhibit superior performance on the two-hop questions than one and three-hop questions, particularly with GPT-3.5 backbones. This is likely because the solution parsing prompts sampled from the SoAyBench are biased towards two-hop questions (refer to Table 2 for more information), causing the GPT-3.5 backbones to default to two-hop solutions within the SoAy framework. On the contrary, GPT-4’s robust instruction-following ability enables it to perform consistently across all question types. Third, GPT-DFSDT exhibits a stronger understanding of one-hop questions when using GPT-3.5 backbones. It breaks down complex questions into simpler one-hop questions, determining the appropriate APIs and their parameters step by step, which enhances its comprehension of one-hop questions.
E.2 Variances
Notably, the data fluctuations marked in Table 3 do not represent variances represented by repeating the identical experiment on identical samples. Given that the $768$ test samples in SoAyBench originate from $44$ template questions, and each template is filled with different entities to create unique question instances, we compute each metric within a group of instantiated questions from the $44$ question templates. We then calculate the mean and variance across groups of different instantiated questions. Our observations indicate that SoAyGPT-4 and SoAyLLaMA delivers a more consistently stable performance across various error types, which benefits from GPT-4’s enhanced generalization capacity across a variety of entities and sufficient and diverse training data from SoAyBench.
E.3 Detailed Error Analysis
Notably, We do not rectify network errors because SoAy aims to apply the system in real-world use, implying that it is essential to test each model’s stability. During experiments, we attempt to request the models up to 3 times. If all 3 attempts fail, the situation is classified as an EE.
E.3.1 ToolLLaMA (7B)
Shows the highest Execution Error (EE) rates across all question types, which might indicate issues with the model’s ability to generate executable programs consistently. It struggles particularly with one-hop questions, which have the lowest Exact Match (EM) rate and high Wrong Solution (WS) errors. Two-hop and three-hop questions reflect similar challenges with WS and EE, although DS errors decrease slightly as the complexity increases. High variability in error rates suggests that the model might benefit from training focused on error reduction techniques. The overall score of 16.72 is the lowest among the models, reflecting significant room for improvement.
E.3.2 GPT-DFSDT
Version 3.5: One-hop questions see extremely high DS rates but lower WS and EE, suggesting issues with solution path diversity but not necessarily with execution. Two-hop and three-hop configurations show a balance of errors across DS, WS, and EE, with a slightly better EM rate for three-hop.
Version 3.5-16K: This model shows improved performance over the standard 3.5 version, with particularly notable improvements in EM and ACC for one-hop questions. DS errors are significantly reduced compared to the 3.5 version, indicating enhanced solution alignment.
Version 4: The most advanced version, showing the best performance with notably high scores, especially in one-hop questions which almost reach perfect EM and ACC. Lower WS rates across all configurations, indicating robustness in generating correct solutions.
E.3.3 SoAyGPT
Version 3.5: Generally strong in EM and ACC, particularly for one-hop questions, suggesting effective solution generation and execution. Lower EE than ToolLLaMA, indicating more reliable program execution.
Version 3.5-16K: Consistently high EM rates across question types, with particularly low DS errors in two-hop questions. Shows good balance across all error types, with no extreme weaknesses.
Version 4: Outstanding performance in one-hop questions with nearly no errors and very high EM and ACC. Two-hop and three-hop questions also show very high scores, making this version highly effective and reliable.
E.3.4 SoAyLLaMA
Chat-7B: Exhibits nearly perfect performance in one-hop questions with zero DS, WS, and WP errors. Three-hop questions maintain high performance but see a slight increase in WP errors, which still remain low compared to other models.
Code-7B: Solid performance with minimal error rates across all question types. A slight uptick in WP errors in two-hop questions could be a focal point for model refinement.
Code-13B: The best-performing model across all metrics, with impressive EM and ACC in one-hop questions and consistent strength in multi-hop configurations. Nearly negligible error rates, setting a high standard for large language models.
Appendix F Other Scenarios
F.1 SoAy implemented on CrossRef
F.1.1 Introduction
Crossref is a registration agency for DOIs (Digital Object Identifiers) established by a consortium of renowned global publishers. It registers DOI numbers for newly published and previously published journal articles. All works registered with Crossref by various publishers automatically establish citation links with each other, facilitating readers to navigate journals and access the websites of different publishers, thus significantly increasing the chances of journal articles being accessed and cited. Crossref’s open reference linking system establishes citation links to journal articles between different publisher websites. For developers, Crossref provides rich open-source API interfaces on GitHub [^6] https://github.com/CrossRef/rest-api-doc?tab=readme-ov-file, such as search and other functionalities.
F.1.2 Implementation
The codes, prompt, and APIs of this part of the implementation are available at our GitHub repo: https://github.com/RUCKBReasoning/SoAy/tree/master/scenarios/CrossRef. As shown above, we need to finish two primary stages: the offline solution library preparation stage and the online solution-based inference stage.
Figure 7: Illustration of Crossref SoAPI Library, SoAPI Graph and solution library.
API Library. As shown in Figure 7, we use 6 APIs of Crossref. The phenomenon of API coupling is prevalent among these SoAPIs. For example, the parameter doi for getWorksByDoi(·) must first be obtained by searchWorkByTItle(·). This necessitates a specific execution order where searchWorkByTItle(·) must be used before getWorksByDoi(·), and a specific execution composition where if getWorksByDoi(·) is used, searchWorkByTItle(·) is definitely used.
Solution Library Construction. The solution library is the cornerstone of the entire framework. Just as what we did for AMiner, we finish the solution library for Crossref, proving SoAy’s usability. The initial stage of SoAy involves analyzing Crossref’s SoAPI and constructing the SoAPI Graph. Following this, a process of sampling and filtering paths is applied to the SoAPI Graph, resulting in the solution library. Based on it, the prompt library is then constructed.
To better understand the performance of Soay on Crossref, let’s take a detailed look at Crossref’s prompts and cases on Figure 10
General Prompt Drafting. The General Prompt outlines a structured approach to crafting programs that utilize a set of APIs to address various queries related to literature and authors. It starts by presenting the APIs that will be employed, such as ‘searchBook‘, ‘getBook‘, and others, which are instrumental in searching for books, obtaining book details, and accessing author information. The prompt then proceeds to illustrate the process through a series of examples. The first example guides the user on how to use the ‘searchBook‘ API to find publishers in a particular subject area (XXX), iterating through the results to compile a list of publisher names. This demonstrates the initial step of narrowing down the scope of the query to a more manageable list. Following that, the second example builds upon the first by not only identifying publishers in a specific field but also by delving deeper to list articles or works associated with those publishers. This is achieved by chaining the ‘searchPublisherBySubject‘ API to ‘getPublisherBasicInfo‘ and then to ‘getPublisherWorks‘, showcasing a more complex use of the APIs to extract detailed information. The prompt concludes with a user query placeholder, encouraging the LLMs to apply the learned methods to formulate a solution. This involves selecting the correct sequence of APIs to retrieve the necessary information in response to the query. The prompt is designed to be a comprehensive guide that adapts to different levels of complexity in the user’s requests, from simple to more advanced searches.
F.1.3 Case Study
Case 1: Mathematics Publishers. The first case presents a query for a list of publishers in the mathematics field. The solution uses the ‘searchPublisherBySubject‘ API with ’math’ as the input to extract a list of publishers, along with the number of DOIs they’ve published. The result provides a collection of publisher names and their respective DOI counts, highlighting their impact in the field of math.
Case 2: DOI Numbers in Computer Science. The second case addresses the request for current DOI numbers from publishers in computer science. By utilizing ‘searchPublisherBySubject‘ to identify relevant publishers, followed by ‘getPublisherBasicInfo‘ to obtain DOI counts, the solution compiles a list of publishers with their current DOI numbers, showcasing their recent publication activities.
Case 3: Article Titles in Computer Science. The third case involves listing articles in the computer science domain. It starts with ‘searchPublisherBySubject‘ to find publishers, then uses ‘getPublisherBasicInfo‘ and ‘getPublisherWorks‘ to retrieve a list of article titles. The output is a concise list of article titles that represent the range of research within computer science.
F.2 SoAy implemented on OpenLibrary
F.2.1 Introduction
Open Library is an online project aimed at creating a comprehensive database of all published books. Operated by the Internet Archive, the website has been serving the public since 2006. It provides information pages for thousands of books, displaying their summaries, covers, publication details, and more. Additionally, it features personal pages for numerous authors, detailing their biographical information, major works, and areas of expertise. For developers, the website also offers web APIs such as search functionality, which are documented in the Public APIs project.
F.2.2 Implemention
The codes, prompt, and APIs of this part of the implementation are also available at our GitHub repo: https://github.com/RUCKBReasoning/SoAy/tree/master/scenarios/OpenLibrary. The same as AMiner and CrossRef, we need to finish two primary stages: the offline solution library preparation stage and the online solution-based inference stage.
Figure 8: Illustration of OpenLibrary SoAPI Library, SoAPI Graph and solution library.
API Library. As shown in Figure 8, we use 6 APIs of Crossref. The phenomenon of API coupling is prevalent among these SoAPIs. For example, the parameter subject for searchSubject(·) must first be obtained by getBook(·) or getAuthorBasicInfo(·). This necessitates a specific execution order where getBook(·) or getAuthorBasicInfo(·) must be used before searchSubject(·), and a specific execution composition where if searchSubject(·) is used, getBook(·) or getAuthorBasicInfo(·) is definitely used.
Solution Library. The solution library is the cornerstone of the entire framework. Just as what we did for AMiner, we finish the solution library for OpenLibrary, proving SoAy’s usability. The initial stage of SoAy involves analyzing OpenLibrary’s SoAPI and constructing the SoAPI Graph. Following this, a process of sampling and filtering paths is applied to the SoAPI Graph, resulting in the solution library. Based on it, the prompt library is then constructed.
General Prompt. To better understand the performance of Soay on OpenLibrary, let’s take a detailed look at OpenLibrary’s prompts and cases on Figure 11 It starts by presenting the APIs that will be employed, such as ‘searchBook‘, ‘getBook‘, and others, which are instrumental in searching for books, obtaining book details, and accessing author information. The prompt then proceeds to illustrate the process through a series of examples. The first example guides the user on how to use the ‘searchBook‘ API to find publishers in a particular subject area (XXX), iterating through the results to compile a list of publisher names. This demonstrates the initial step of narrowing down the scope of the query to a more manageable list. The same principle applies to the two other examples. The prompt concludes with a user query placeholder, encouraging the LLMs to apply the learned methods to formulate a solution. This involves selecting the correct sequence of APIs to retrieve the necessary information in response to the query. The prompt is designed to be a comprehensive guide that adapts to different levels of complexity in the user’s requests, from simple to more advanced searches.
F.2.3 Case Study
Case 1: Identifying Mathematics Publishers. In the first case, the objective is to identify publishers that specialize in mathematics. The process begins by utilizing the ‘searchPublisherBySubject‘ API, specifying ’math’ as the subject parameter. This call returns a list of publishers active in the field of mathematics, each accompanied by the number of DOIs (Digital Object Identifiers) they have generated. The output is a curated list of publisher names alongside their DOI counts, which serves as an indicator of their contributions and presence in the mathematical literature space.
Case 2: Obtaining DOI Numbers for Computer Science Publishers. The second case focuses on obtaining the current DOI numbers for publishers within the computer science field. The ‘searchPublisherBySubject‘ API is employed to first filter out publishers related to computer science. Once a list of these publishers is obtained, the ‘getPublisherBasicInfo‘ API is then used to fetch the up-to-date DOI numbers for each publisher. The result is a compilation of publisher names paired with their current DOI counts, providing a snapshot of their recent scholarly output and engagement within the field.
Case 3: Listing Computer Science Articles. The third case aims to extract a list of article titles from the computer science domain. The process initiates with the ‘searchPublisherBySubject‘ API to pinpoint publishers in the field, followed by the use of ‘getPublisherBasicInfo‘ to acquire the publisher’s unique ID. With the publisher ID in hand, the ‘getPublisherWorks‘ API is then applied to retrieve a comprehensive list of article titles associated with the publisher. The final output is a succinct array of article titles, offering a glimpse into the diverse range of research topics and findings published within the realm of computer science.
Part III. Prompts
In this section, we will comprehensively present the prompts used in the data generation process of SoAy, as well as the implementation process of SoAyGPT on AMiner and two other scenarios.
Figure 9: A Preview of the prompt library. $P_{s}$ signifies the prompt for solution retrieval, $P_{p}$ represents the prompt for program generation, and $P_{a}$ denotes the prompt for answer generation. All the black text except the ellipsis is the original prompt content, the yellow highlighted text is the annotations added for better understanding in the context of creating a prompt, and the purple highlighted text is the content waiting for the LLM output.
Figure 10: A Preview of the prompt library and three cases for Crossref.
Figure 11: A Preview of the prompt library and three cases for OpenLibrary.
Figure 12: A Preview of the prompts used in the data preparation process.
Figure 13: The screenshots of the demonstration systems https://soay.aminer.cn and online service widget https://chatglm.cn/main/gdetail/65bf5a99396389a73ace9352 powered by SoAy.

### v2

@@
-for Academic Information Seeking
-Abstract
-Applying large language models (LLMs) for academic API usage shows promise in reducing researchers’ academic information seeking efforts.
-However, current LLM API-using methods struggle with complex API coupling commonly encountered in academic queries.
-To address this, we introduce SoAy, a solution-based LLM API-using methodology for academic information seeking.
-It uses code with a solution as the reasoning method, where a solution is a pre-constructed API calling sequence.
-The addition of the solution reduces the difficulty for the model to understand the complex relationships between APIs.
-Code improves the efficiency of reasoning.
+SoAy: A Solution-based LLM API-using Methodology for Academic Information Seeking
+Abstract.
+Applying large language models (LLMs) to academic API usage shows promise in reducing researchers’ efforts to seek academic information.
+However, current LLM methods for using APIs struggle with the complex API coupling commonly encountered in academic queries.
+To address this, we introduce SoAy, a solution-based LLM methodology for academic information seeking.
+SoAy enables LLMs to generate code for invoking APIs, guided by a pre-constructed API calling sequence referred to as a solution.
+This solution simplifies the model’s understanding of complex API relationships, while the generated code enhances reasoning efficiency.
+LLMs are aligned with this solution-oriented, code-based reasoning method by automatically enumerating valid API coupling sequences and transforming them into queries and executable code.
 To evaluate SoAy, we introduce SoAyBench, an evaluation benchmark accompanied by SoAyEval, built upon a cloned environment of APIs from AMiner.
 Experimental results demonstrate a 34.58-75.99% performance improvement compared to state-of-the-art LLM API-based baselines.
-All datasets, codes, tuned models, and deployed online services are publicly accessible at https://github.com/RUCKBReasoning/SoAy.
-[Uncaptioned image] A Solution-based LLM API-using Methodology
-for Academic Information Seeking
-Yuanchun Wang, Jifan Yu§ [^1] Work was done when these authors interned at Zhipu AI., Zijun Yao§, Jing Zhang, Yuyang Xie§, Shangqing Tu§ Yiyang Fu, Youhe Feng, Jinkai Zhang, Jingyao Zhang, Bowen Huang, Yuanyao Li Huihui Yuan, Lei Hou§, Juanzi Li§and Jie Tang§ Renmin University of China      §Tsinghua University      Zhipu AI https://soay.aminer.cn
-1 Introduction
+All codes are publicly accessible at https://github.com/RUCKBReasoning/SoAy.
+Large Language Model, API, Academic Information Seeking
+journalyear: 2025copyright: acmlicensedconference: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 3–7, 2025; Toronto, ON, Canadabooktitle: Proceedings of the 31st ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’25), August 3–7, 2025, Toronto, ON, Canadadoi: 10.1145/3690624.3709412isbn: 979-8-4007-1245-6/25/08ccs: Computing methodologies Multi-agent planningccs: Computing methodologies Semantic networksccs: Information systems Data mining
+1. Introduction
 Researchers frequently seek academic information such as a paper’s metadata, a scholar’s publication record, and the connections between scholars as part of their routine scholarly activities.
-The advent of academic information retrieval systems—including DBLP dblp team (2005), CiteSeer Giles et al.
-(1998), and Google Scholar Noruzi (2005)—and more sophisticated academic mining platforms such as Microsoft Academic Search Sinha et al.
-(2015) and AMiner Tang et al.
-(2008), has substantially reduced the complexity of exploring the academic data, which encompasses massive entities like authors, conferences, and papers.
-Figure 1: Comparing responses from Chameleon Lu et al. (2023), ToolLLaMA Qin et al. (2023b), and the proposed SoAy, all leveraging AMiner APIs to address the same academic query.
-Large Language Models Zhao et al.
-(2023) (LLMs) unveil the potential to further reduce the effort required for academic information seeking, benefiting from LLMs’ reasoning ability to comprehend intentions expressed in natural language.
-Enabling LLMs to utilize existing academic retrieval/mining APIs Qin et al.
-(2023a) tends to be a feasible solution.
-However, this adaptation is not easy because of two main challenges:
-(1) Coupling: Academic information for academic queries is distributed across various storage locations but interconnected internally, resulting in multiple academic APIs responsible for different functions yet interconnected, i.e.
-coupled.
-This requires LLM to fully understand the sequential order of calls between these APIs.
-Existing API-using methods like Chameleon Lu et al.
-(2023), shown in Figure 1(a), retrieve and execute APIs separately to get results, thus failing to handle such API coupling.
-(2) Efficiency: Existing methods such as ToolLLM Qin et al.
-(2023b) illustrated in Figure 1(b) employ a depth-first search-based decision tree (DFSDT) for step-by-step reasoning, which has the potential to capture the regularities in the sequence of API.
-However, the excessive steps lead to query times that users may find unacceptable, compared with traditional academic search engines that offer researchers a very fast response.
-We propose SoAy, a Solution-based API-using methodology to apply LLMs to for academic information seeking.
-Our key innovation lies in the new API calling method: solution and code.
+The advent of academic information retrieval systems—including DBLP (dblp team, 2005), CiteSeer (Giles et al., 1998), and Google Scholar (Noruzi, 2005)—and more sophisticated academic mining platforms such as Microsoft Academic Search (Sinha et al., 2015) and AMiner (Tang et al., 2008), has substantially reduced the complexity of exploring the academic data, which encompasses massive entities like authors, conferences, and papers.
+However, existing systems are constrained by their predefined functionalities and may not fully address the flexible and complex needs of users.
+For example, when a researcher aims to know how many times the representative work of Yann LeCun has been cited, he or she often handcrafts the search keywords into Google Scholar search box, finds and clicks the target item on the result page, and finally seeks the target publication and its citation count.
+Large Language Models (LLMs) (Zhao et al., 2023), with their robust reasoning capabilities, offer a promising solution for comprehensively understanding and addressing such nuanced requirements.
+One potential approach involves enabling LLMs to leverage academic seeking APIs (Qin et al., 2023a).
+The diverse functionalities offered by these APIs, when combined effectively, hold the potential to address a wide array of user needs.
+However, this adaptation presents two significant challenges:
+(1) Coupling: One key challenge in adapting large language models (LLMs) to leverage academic retrieval/mining APIs is the inherent parameter coupling within these APIs.
+For instance, to find the most cited paper of a scholar, the process involves identifying the scholar, retrieving their papers, and then sorting by citation count.
+This sequence requires the LLM to understand the interdependent calls between the scholar and publication search APIs (i.e., searchPerson and getPersonPubs APIs).
+Existing methods, such as Chameleon (Lu et al., 2023), as depicted in Figure 1(a), retrieve and execute APIs independently, overlooking the necessary coupling.
+Consequently, they fail to correctly invoke the getPersonPubs API, which requires a person ID provided by the searchPerson API.
+(2) Efficiency: Existing methods, such as ToolLLM (Qin et al., 2023b), employ a depth-first search-based decision tree (DFSDT) for step-by-step reasoning, which can effectively capture the couplings in the sequence of API calls.
+However, the numerous steps involved result in multiple inference times, which may be unacceptable to users who are accustomed to the rapid response of existing academic information retrieval systems.
+For instance, as shown in Figure 1(b), to obtain the most cited paper of a scholar, DFSDT requires one LLM inference per returned paper from the getPersonPubs API to determine its citation count, leading to high inference costs.
+Figure 1. Comparing responses from Chameleon (Lu et al., 2023), ToolLLaMA (Qin et al., 2023b), and the proposed SoAy, all leveraging AMiner APIs to address the same academic query.
+[…]
+Comparing responses from Chameleon (Lu et al., 2023), ToolLLaMA (Qin et al., 2023b), and the proposed SoAy, all leveraging AMiner APIs to address the same academic query.
+Figure 2.
+The overall SoAy process involves a single academic question case.
+The main steps of data generation are illustrated on the left and middle sections.
+Two modes of LLM alignment with the automated triplet dataset are on the right.
+[…]
+The overall SoAy process involves a single academic question case.
+The main steps of data generation are illustrated on the left and middle sections.
+Two modes of LLM alignment with the automated triplet dataset are on the right.
+We propose SoAy, a Solution-based API-using methodology to apply LLMs for academic information seeking.
+Our key innovation is the solution-oriented, code-based method for API calling.
 As shown in Figure 1(c), SoAy first lets the LLM generate a feasible API calling plan, i.e.
-solution, based on complex user inputs, and then allows the LLM to generate executable API calling code based on the generated solution.
-The pre-generated solutions can assist the LLM in clarifying the complex coupling relationships between academic APIs when generating subsequent code.
-Calling APIs through code generation in a single inference process without repeated reasoning can achieve efficiency levels closer to traditional academic search systems.
-Specifically, the ability for LLM to reliably utilize complex academic APIs in this manner to accomplish academic search tasks benefits from the automatic generation of a training dataset, consisting of (query, solution, code) triplets.
-The data is created on the AMiner APIs, resulting in 3,960 automated triplets.
-We retain 4/5 of these triplets, providing valuable resources for fine-tuning existing LLMs or demonstrating LLM in-context learning.
-This alignment helps bridge the gap between general LLMs and the intricacies of academic API usage.
-To evaluate SoAy, we utilize 1/5 of the generated data, comprising 792 (query, solution, code) triplets.
-These are enhanced through human verification to ensure the quality of the test set.
+solution, based on complex user inputs, and then allows the LLM to generate executable API calling code based on this solution.
+The pre-generated solutions help the LLM clarify complex coupling relationships between academic APIs when generating subsequent code.
+By generating the API calling code in a single inference process, SoAy achieves efficient inference without repeated reasoning.
+To enable LLMs to reliably generate solutions and code in this manner, it is crucial to automatically create a training dataset consisting of (Query, Solution, Code) triplets.
+The process involves using academic APIs published by aminer.org to create a solution library by constructing a dependency graph among the APIs and identifying valid API calling paths.
+Each solution is then linked with input parameters and output attributes to align more precisely with user intentions.
+This binding is transformed into a template query and then into Python code.
+Code correctness is ensured by instantiating the input, executing the code, and verifying the results against expected outputs.
+As a result, we automatically develop a dataset containing 3,960 (query, solution, code) triplets, consisting of SoAyBench.
 We then introduce SoAyEval, an evaluation methodology that takes into account both solution correctness and code execution accuracy.
-By employing SoAyEval for automated evaluation alongside human assessment, our experiments demonstrate the effectiveness and efficiency of SoAy in utilizing AMiner APIs compared to state-of-the-art LLM API-using baselines, across both in-context learning and tuning modes.
+The experiments demonstrate the effectiveness and efficiency of SoAy in utilizing AMiner APIs compared to state-of-the-art LLM API-using baselines, across both in-context learning and tuning modes.
 Furthermore, we validate the versatility of SoAy for application in other academic retrieval systems.
-The paper contributes: (1) SoAy, a novel LLM adaptation method, addressing challenges like coupling and efficiency in academic information seeking through automatic data generation, teaching LLMs to use academic APIs.
+The paper contributes: (1) SoAy, A novel LLM adaptation method that addresses challenges such as coupling and efficiency in academic information seeking.
+This method involves automatic data generation, enumerating API calling sequences, transforming them into queries and executable code, and teaching LLMs to use academic APIs.
 (2) SoAyBench, a released benchmark dataset, accompanied by the SoAyEval evaluation method, to assess LLMs’ proficiency in using academic APIs.
-(3) Demonstrated effectiveness and efficiency of SoAy in the AMiner system, along with its real deployment on an online service as shown in Figure 13.
-Figure 2: Illustration of AMiner API Library, API Graph and solution library.
-2 Task Formulation
-The task for SoAy is to seek appropriate academic information $I$ based on complicated user query $Q$ in an efficient way.
-This objective tends to be achieved by using academic APIs in an valid and efficient order.
-Problem 1
-User Intent Parse Based on $Q$, SoAy needs to understand which academic information the user requires, then identify the associations within them and conduct separate queries.
-$Q\rightarrow Q_{S},\quad Q_{S}=\{Q_{i}\}_{i=1}^{N}.$ (1)
-Traditional academic search systems cannot accomplish this task, thus requiring users to break their requirements $Q$ down into step-by-step queries $Q_{S}$ to type in the search box.
-SoAy, on the other hand, needs to use LLMs to break the intent down and get the relevant API calling sequence.
-Problem 2
-Academic Information Aggregation After understanding what information the user truly needs, SoAy’s task is to retrieve relevant information from data sources according to $Q_{S}$ and aggregate and process the results to provide the user with the precise information $I$ they seek.
-Traditional information retrieval systems lack this design and instead display all retrieved information to users for them to select what they want in a list-like format.
-What SoAy needs to do is generate the correct code to call the academic APIs and process the obtained information within a comparable time to those traditional systems.
-Figure 3: The overall process of SoAy around one case of an academic question.
-The main steps of data generation are shown on the left.
-Two mode of LLM alignment with the automated triplet dataset are on the right.
-3 SoAy: Solution-based API-using
-We introduce SoAy to adapt general LLMs for academic information seeking, emphasizing the use of existing academic APIs.
-To achieve this, abundant training data in the form of (academic query, executable API sequence) pairs is essential, facilitating either fine-tuning of existing LLMs or guiding them through in-context learning.
-Illustrated in the left part of Figure 3, this data generation pipeline comprises four key steps:
-(1) Solution Library Construction.
-Utilizing the provided academic APIs, we construct a graph representing the dependencies among APIs, and traverse this graph to identify valid and simple paths, referred to as solutions.
+(3) Demonstrated effectiveness and efficiency of SoAy in the AMiner system, along with its real deployment on an online service.
+2. SoAy: Solution-based API-using
+2.1. Background
+The process users seek academic information using traditional academic information retrieval systems can be summarized as follow:
+Intent Decomposition: To address their complex information needs, users often begin by formulating keywords and entering them into a search interface.
+For example, to seek the answer of the question illustrated in Figure 1, a user might input “New York University Yann Lecun” into the search box of the academic information retrieval system to explore relevant results.
+Step-by-step Action: Users then follow a step-by-step process to find the information they need.
+This involves searching in the search box, browsing through the retrieved scholars, selecting the target scholar, and exploring further by continued browsing and clicking.
+This sequence of user actions essentially corresponds to a series of API calls that underpin traditional academic systems.
+For instance, entering a query in the search box translates to calling the searchPerson API as depicted in Figure 3.
+Clicking on a scholar in the search results corresponds to calling APIs like getPersonPubs to retrieve more detailed scholar information.
+Finally, clicking on a paper with the most citations to find out its citation count involves calling the getPublication API.
+Thus, in this scenario, the API calling sequence is: searchPerson $\rightarrow$ getPersonPubs $\rightarrow$ getPublication.
+This sequence of API calls is defined as a solution, as detailed in Definition 3.2.
+Figure 3. Illustration of AMiner API Library $L_{A}$, API Graph $G$ and solution library $L_{S}$.
+[…]
+Illustration of AMiner API Library $L_{A}$, API Graph $G$ and solution library $L_{S}$.
+2.2. SoAy Framework
+To tailor general LLMs to comprehend user requirements and methodically invoke academic APIs correctly and efficiently to access the desired information, we utilize pre-collected solutions $S$ to aid the LLM in crafting Python code that incorporates the necessary API calls.
+To achieve this, abundant training data in the form of (Query, Solution, Code) triplet is essential, facilitating either fine-tuning of existing LLMs or guiding them through in-context learning.
+Illustrated in the left and middle parts of Figure 2, this solution collection and data generation pipeline comprises four key steps:
+(1) Solution Library Construction (Section 3).
+Utilizing the provided academic APIs, we construct a graph representing the dependencies among APIs, and traverse this graph to identify valid and concise paths, referred to as solutions.
 These solutions collectively form a library known as the solution library.
-(2) Combination Formulation.
-To obtain a more precise semantic of each solution in the library, we combine all the possible inputs of the initial API and all the outputs of the final API of the solution to formulate a series of combinations.
-(3) Query Generation.
+(2) Combination Formulation (Section 4.1).
+To obtain more precise semantics of each solution in the library, we combine all the possible inputs of the initial API and all the outputs of the final API of the solution to formulate a series of combinations.
+(3) Query Generation (Section 4.2).
 We employ ChatGPT to transform each combination into a template question lacking specific academic entity data.
-These templates are then diversified into various expressions and populated into potential academic queries with data from a specialized academic source, the AMiner database here.
-(4) Code Generation.
-Finally, we use ChatGPT again to write a piece of corresponding code given the combination, which is expected to correspond to the template query that shares the same combination.
-Lastly, we employ ChatGPT once more to generate the corresponding code for the given combination.
-This code is anticipated to match the template query that has the same combination.
-Notably, in the data preparation part, we indicate ChatGPT as gpt-3.5-turbo-0613. We present the details of the data preparation prompt and the process of setting up SoAyBench in the appendix.
-4 Solution Library Construction
+These templates are then diversified into various expressions and transformed into specific academic queries, populated with concrete academic entities as input and output from the specialized academic source, the AMiner database.
+(4) Code Generation (Section 4.3).
+Finally, we use ChatGPT to generate Python code corresponding to the given combination, which is expected to handle various queries that share the same combination.
+For each instantiated query, we ensure code correctness by incorporating the query input into the code, executing the code, and verifying the results against the query’s ground truth outputs.
+Notably, during the process of data creation, we indicate ChatGPT as gpt-3.5-turbo-0613.
+3. Solution Library Construction
 Given an academic API Library $L_{A}$, the initial step is to build an API graph where each node represents an API $f$, and a directed edge from node $f_{n}$ to node $f_{m}$ signifies a coupling relationship from $f_{n}$ to $f_{m}$ with the intersected attributes $a_{n,m}$.
-Definition 1
+Definition 0.
 API Coupling defines the frequent dependencies of the inputs and outputs in academic API calling, caused by academic entity correlations’ complexity:
-$f_{n}$ $\xrightarrow{a_{n,m}}f_{m},$ (2)
-$if\quad a_{n,m}=i_{m}$ $\cap o_{n}\quad\text{and}\quad a\neq\emptyset,$
-where $i_{n}$ denotes the input parameters of $f_{m}$ and $o_{n}$ denotes the returned attributes of $f_{n}$.
-[…]
-Definition 2
-A Solution $S$ is defined as an execution sequence of APIs traversed from the API graph:
+(1) $f_{n}$ $\xrightarrow{a_{n,m}}f_{m},$
+$if\quad a_{n,m}=i_{m}$ $\cap o_{n}\quad\text{and}\quad a_{n,m}\neq\emptyset,$
+where $i_{m}$ denotes the input parameters of $f_{m}$ and $o_{n}$ denotes the returned attributes of $f_{n}$.
+The API coupling $f_{n}\xrightarrow{a_{n,m}}f_{m}$ means that we can use $f_{n}$ to output the attribute $a_{n,m}$, and subsequently use $a_{n,m}$ as the input parameter to call $f_{m}$.
+This is because $a_{n,m}$ is the intersection of the returned attributes $o_{n}$ of $f_{n}$ and the input parameters $i_{m}$ of $f_{m}$.
+The graph $G$ built from AMiner APIs according to their coupling relationships is shown in Figure 3(b).
+By traversing this graph while limiting the maximum number of hops to $H$, we can derive an API calling sequence, referred to as a solution:
+Definition 0.
+A Solution $S$ is defined as a calling sequence of APIs traversed from the API graph:
 $S=f_{1}\xrightarrow{a_{1,2}}f_{2}\cdots\xrightarrow{a_{k-1,k}}$ $f_{k}\cdots\xrightarrow{a_{J-1,J}}f_{J},$
-$\forall k,f_{k}\in$ $L_{A}.$ (3)
+(2) $\forall k,f_{k}\in$ $L_{A}.$
 A solution library $L_{S}$ denotes a set of valid and simple solutions, ensuring by the following ways.
-Valid Solution. Among all traversed solutions from the API graph, validity is determined by starting with a fuzzy API, as only fuzzy APIs can accept user utterances in natural language as input.
-Simple Solution.
+Valid Solution.
+A valid solution can be checked by judging whether the node in the API graph corresponding to the first API in this solution has zero Indegree.
+Note this is specific to academic information seeking systems.
+The situation may vary for other application systems.
+Concise Solution.
 For a single query, there may be multiple solutions available.
-The preferred choice is the simplest one, containing the fewest APIs.
-For instance, when inquiring about the publication date of the representative work of a scholar, the optimal solution is “searchPerson $\rightarrow$ getPersonPubs”, despite obtaining the same information is also possible through “searchPerson $\rightarrow$ getPersonPubs $\rightarrow$ getPublication $\rightarrow$ searchPublication”.
-We meticulously craft prompts to instruct ChatGPT in fulfilling the above two characteristics.
-All prompts for question generation are detailed in Appendix Part III.
-Prompts.
+The preferred choice is the most concise one, containing the fewest APIs.
+We utilize ChatGPT to verify if two given solutions can obtain the same information.
+The specific prompt used for this verification can be found in our code repo.
+Specifically, each traversed solution is sequentially added to the solution library $L_{S}$.
+For each new solution $S$, we check whether a more concise solution already exists in $L_{S}$ that can achieve the same result as $S$.
 Consequently, we’ve created 16 solutions to establish the solution library $L_{S}$ for AMiner APIs.
 Figure 3(c) presents a subset of these solutions, including their input parameters, returned attributes, and corresponding question templates.
-5 Reasoning Code Generation
-In order for LLM to aggregate effective academic information and provide users with precise answers, SoAy let LLM generate code containing API calls for academic information retrieval.
-To achieve this, a large amount of Query, Corresponding Code data needs to be generated to train LLM.
-5.1 Combination Formulation
-Section 2 describes the one-to-several relationship between a query and multiple solutions.
-It’s evident that a similar one-to-several relationship exists between a solution and several query intents.
-In a solution $S$, the first API is referred to as the head API $f_{1}$, while the last API is termed the tail API $f_{J}$.
-For example, in Figure 2(c), consider the 3-hop solution “searchPublication $\rightarrow$ getPublication $\rightarrow$ getPersonBasicInfo”.
-Due to the diversity of the Head API Input $i_{1}$ and the Tail API Output $o_{J}$, determining which information (outputs) is sought under what conditions (inputs) remains uncertain even after the solution is established.
-An intuitive approach is to combine the head API input, the solution, and the tail API output to refine the intent space into a specific requirement.
-A solution-based combination is defined as a specific binding of the head API input, the solution, and the tail API output: $C=(i_{1},S,o_{J})$.
-Given the $i_{1}$ is “title” and the $o_{J}$ is “education_experience”, the $S$ “searchPublication $\rightarrow$ getPublication $\rightarrow$ getPersonBasicInfo” could lead to one plausible interpretation: querying the graduation school of the first author of the provided publication.
-It’s important to recognize that the information from the academic data source corresponding to $i_{1}$ is both the information needed by q-t to fill and the input information needed by the generated code, $o_{J}$ is both the expected answer of q-t and the golden result of the code.
-5.2 Query Generation
+4. Soluion-based Data Creation
+For each solution in the solution library, $L_{S}$, we generate corresponding queries and code. This enables LLMs to effectively perform academic information seeking.
+4.1. Combination Formulation
+Typically, a solution may correspond to multiple query intents, each representing a distinct need.
+Therefore, we formulate combinations derived from a solution to clearly represent each specific intent.
+We here define a head API $f_{1}$ as the first API in a solution $S$, while the last API is termed the tail API $f_{J}$.
+For example, in Figure 3(c), considering the 3-hop solution “searchPublication $\rightarrow$ getPublication $\rightarrow$ getPersonBasicInfo”, searchPublication is $f_{1}$ and $f_{J}$ is getPersonBasicInfo.
+Due to multiple choices of the Head API Input $i_{1}$ and the Tail API Output $o_{J}$, determining which information (outputs) is sought under what conditions (inputs) remains uncertain even after the solution is established.
+We define a solution-based combination as a specific binding of the head API input, the solution, and the tail API output: $C=(i_{1},S,o_{J})$, to refine the solution $S$ into a targeted intent.
+For example, given $i_{1}$ is “title” and $o_{J}$ is “education_experience”, the solution “searchPublication $\rightarrow$ getPublication $\rightarrow$ getPersonBasicInfo” can refer to a specific interpretation: querying the alma mater of the first author of the provided publication.
+4.2. Query Generation
 Template-Query Generation.
-For a given combination $C$, we define a template question $Q_{T}$ to represent the intent of $C$ with missing entity information.
-In the example provided, the corresponding template question $Q_{T}$ could be formulated as “Which school did the first author of XXXX graduate from?”, where “XXXX” refers to a placeholder representing the title of a publication, as illustrated in Table 1.
-Semantic Expansion and Entity Filling.
-A specific intent represented by a template question $Q_{T}$ may be expressed by multiple natural language queries.
-For instance, “Which school did XXX graduate from?” and “What is XXX’s alma mater?” convey the same intent.
-To enrich the semantic diversity of an intention, we conduct semantic augmentation by crafting prompts to leverage ChatGPT in generating three new, semantically enriched template questions for each combination.
-Following semantic augmentation, we replace the special symbols in the templates (represented by uppercase “X”s) with information about a specific academic entity in an academic data soruce.
-A placeholder in a question template must be replaced by the attributes of specific academic entities, such as person_id, name, interest, organization, bio, educational_experience, etc., as illustrated in Figure 2.
-Taking the instance of the alma mater query again, The name of this scholar will be used to fill in the $Q_{T}$ to become a real question “What is scholar name’s alma mater” and the educational experience of this scholar serves as the golden answer to this query.
-Table 1: The semantics of symbols in $Q_{T}$.
+For each combination $C$, we define a template question $Q_{T}$ to capture its intent.
+For instance, for the aforementioned combination C=C=(“title”, “searchPublication $\rightarrow$ getPublication $\rightarrow$ getPersonBasicInfo”, “education_experience”), the template question $Q_{T}$ could be phrased as, “Which school did the first author of XXXX graduate from?” where “XXXX” serves as a placeholder for the “title of a publication”, as shown in Table 1.
+Additionally, $Q_{T}$ could be expressed in variations like, “What is the alma mater of the first author of XXXX?” to convey the same intent.
+To broaden the semantic diversity of an intention, we engage in semantic augmentation by prompting ChatGPT to generate three varied template questions for each combination, enhancing the range of phrasing and nuances.
+Template-Query Instantiation.
+For each template query $Q_{T}$, we substitute the placeholders denoted by uppercase “X”s with specific attributes such as person_id, name, interest, organization, bio, and educational_experience of academic entities from an academic data source to create the instantiated query $Q$.
+For instance, using the alma mater query as an example again, the placeholder in $Q_{T}$ is replaced with the actual name of a scholar denoted by $i_{1}^{\prime}$, resulting in a specific question like “What is scholar name’s alma mater?”.
+The actual educational background of this scholar denoted by $o_{j}^{\prime}$, then serves as the ground truth answer to $Q$.
+Table 1. The semantics of symbols in $Q_{T}$.
 Symbol Semantics
 X interest
-XX organization
+$X$ organization
 XXX name
@@
 Y Year
-5.3 Code Generation
-In the final step, we employ ChatGPT again to generate a piece of code that corresponds to a given combination $C$ with missing input information.
-This code is intended to get the answer of the template query $Q_{T}$ shared the same combination $C$ with it.
-Still with information from a specific academic entity, the $i_{1}$ serves as the input of the generated code.
-The code is expected to be executed to get the same result as $o_{J}$, which can be seen as a golden information of this code.
-If the result matches the golden information, this code is valid.
-Any invalid code, not producing matching results, will be excluded from the final dataset.
-6 LLMs Alignment
-In order to ultimately leverage LLM to accomplish the task of academic information seeking, SoAy needs to align LLM to be capable of calling academic APIs correctly and efficiently through generating solutions and code.
-This can be achieved by constructing training data and aligning the LLMs according to it, whether through a tuning or tuning-free method.
-6.1 SoAyBench–{(Query, Solution, Code)}
-Table 2: Question statistics in SoAyBench set.
-Question Type One-hop Two-hop Three-hop Total
+4.3. Code Generation
+So far, a solution $S$ is tailored to a specific intent by associating head API input $i_{1}$ and tail API output $o_{J}$ to form a combination $C=(i_{1},S,o_{J})$, which is linked to three template queries $\{Q_{T}\}$.
+Each $Q_{T}$ is instantiated with real academic entities, resulting in a set of specific queries $\{Q\}$, each with its own input $i_{1}^{\prime}$ and corresponding ground truth $o_{J}^{\prime}$.
+Building on this data, we then utilize ChatGPT to generate a piece of code corresponding to each combination $C$.
+This code is designed to retrieve answers for various queries that share the same combination.
+Using $i_{1}^{\prime}$ from a query $Q$ as the input for the code, we execute the code and then compare its output to the ground truth $o_{J}^{\prime}$.
+If the output aligns with the ground truth, the code is considered valid.
+Any code that fails to produce matching results is excluded from the final dataset.
+5. LLMs Alignment
+To align a general-purpose LLM with the mode of generating corresponding solutions and code based on given problems, we need to prepare relative training data (Section 5.1) and choose the aligning method (Section 5.2).
+5.1. SoAyBench–{(Query, Solution, Code)}
+Table 2. Query statistics of the Created Dataset.
+Query Type One-hop Two-hop Three-hop Total
 Scholar 540 1,800 540 2,980
@@
 Total 720 1,980 1,230 3,960
-After going through the aforementioned four stages, a substantial amount of high-quality (Query, Solution, Code) data has been automatically collected.
-We organize this data into a comprehensive benchmark known as SoAyBench.
-Currently, SoAyBench consists of 3,960 triplets, out of which one-fifth has undergone expert inspection and adjustment to form a test set.
-This test set is used to evaluate the ability of the LLMs to search academic information using the AMiner API.
-The remaining portion of the data is utilized to support the alignment of the LLMs.
-Table 2 shows the question statistics of SoAyBench.
-Table 3: Overall Performance comparison of the proposed SoAy and the baseline methods ToolLLaMA and GTP-DFSDT, as tested on SoAyBench.
+After going through the aforementioned four stages, a substantial amount of high-quality (Query, Solution, Code) data has been automatically created.
+Table 2 shows the query statistics of the dataset.
+Currently, SoAyBench consists of 17 solutions with 44 corresponding combinations.
+Each combination’s template query is augmented 3 times and then, through 30 times of query instantiation, results in 90 real queries.
+This process yields a total of $44\times 3\times 30=3,960$ triplets.
+One-fifth of these triplets have undergone expert inspection and adjustment to form a test set, which is published as SoAyBench.
+This test set is supported by a publicly available academic API service, cloned from the online AMiner service.
+It is used to evaluate the ability of LLMs to search academic information using the AMiner API.
+The remaining data is used as the training data to support the alignment of LLMs.
+Table 3.
+Overall Performance comparison of the proposed SoAy and the baseline methods ToolLLaMA and GTP-DFSDT, as tested on SoAyBench.
 The terms 3.5, 3.5-16K, and 4 are abbreviations for different versions of GPT.
@@
 […]
-DS(%) WS(%) WP(%) EE(%)
+DS(%) WS(%) WC(%) EE(%)
 ToolLLaMA 7B one-hop 12.50±8.00 24.31±13.26 1.39±0.00 54.17±16.01 7.64±5.20 20.14 16.72
@@
 three-hop 0.00±0.00 0.00±0.00 2.38±2.86 0.40±0.00 97.22±4.28 97.22
-6.2 Model Alignment
+5.2. Model Alignment
 We provide alignment strategy for both super large closed-source LLMs and tunable open-source LLMs.
-(1) SoAyGPT is our proposed solution for models like GPT Eloundou et al.
-(2023), which cannot have their parameters modified but possess strong instruction following capabilities.
-We select a diverse set of triplets from SoAyBench data that represent different usages of solutions as examples to construct prompts.
-Given a query, SoAyGPT generates the solution and code in sequence through in-context learning, followed by executing the generated code to obtain the result.
-We utilize different versions of GPT models as backbones for this purpose.
-(2) SoAyLLaMA, on the other hand, is our alignment strategy for the LLaMA Touvron et al.
-(2023) model, which is an open-source model that we can fine-tune within an acceptable training cost.
-By using the query of the triplets from SoAyBench as input and integrating the solution and code as output, SoAyLLaMA is trained to generate the solution and code in a single step based on the given query in a zero-shot setting.
-7 Experiments
+Closed-source LLMs Alignment.
+SoAyGPT is our proposed solution tailored for models like GPT, which, while unable to have their parameters modified, exhibit strong instruction-following capabilities.
+We engage these closed-source LLMs through in-context learning, leveraging training data from our created dataset to generate prompts.
+We primarily create three types of prompts: solution, code, and answer generation prompts.
+The solution generation prompt is constructed by integrating all solutions explained by their corresponding template queries from the training data.
+For each solution in the solution library $L_{A}$, we create a specific code generation prompt by incorporating various instantiated queries and code snippets corresponding to the solution from the training data.
+Since LLMs may generate solutions not existing in $L_{A}$, we also create a general code generation prompt for any new solution by including instantiated queries and code snippets of all solutions in $L_{A}$.
+The answer generation prompt includes examples of queries, code execution results, and corresponding natural language answers to curate human-friendly responses.
+Upon receiving a query, SoAyGPT first generates a solution using the solution generation prompt.
+If the solution can be found in $L_{S}$, the specific code generation prompt for it is used to generate the code; otherwise, the general code generation prompt is used.
+The generated code is then executed to produce the result.
+Finally, the answer generation prompt is used to compose the query and the execution result into a natural language answer.
+All these prompts are illustrated in Figure 7 in the Appendix.
+Open-course LLMs Alignment.
+SoAyLLaMA, on the other hand, is our alignment strategy for LLaMA (Touvron et al., 2023), which is an open-source model that we can fine-tune within an acceptable training cost.
+By using the instantiated query of the triplets from the training data as input and integrating the solution and code as output, SoAyLLaMA is trained to generate the solution and code in a single inference step based on the given query in a zero-shot setting.
+6. Experiments
 In this section, we conduct experiments to validate the effectiveness, efficiency, and practicality of SoAy.
-We focus on the following Research Questions: (RQ1) In which part of the academic information seeking do LLMs struggle the most?
-(RQ2) In terms of effectiveness on academic information seeking, does SoAy outperform existing methods?
-(RQ3) In terms of efficiency, does SoAy finish the task within an execution time that users can accept?
-(RQ4) What kind of base model is more suitable for constructing SoAy?
-We conduct experiments on SoAyBench to compare SoAy with other baseline methods, focusing on these research questions.
-Additionally, we introduced a detailed evaluation method called SoAyEval, which could also be used in the evaluation of other tool-using scenarios.
-We also perform online evaluations specifically targeting real-world academic information seeking use cases.
-There’s an independent chapter in the appendix showing more implementation details and interesting findings.
-7.1 Experimental Settings
-Baselines.
-Many existing tool-utilizing works fail to function effectively due to API inter-dependency and intricate execution logic.
-We have chosen GPT-DFSDT, employing the same three versions of the GPT as backbone as SoAyGPT and ToolLLaMA-7b Qin et al.
-(2023b) as baselines in API-using experiment on SoAyBench, as the former represents the state-of-the-art model on a general domain tool-utilizing benchmark, ToolBench Qin et al.
-(2023b), and the latter is an open-source model that has been fine-tuned on numerous tool-utilizing data samples, achieving results comparable to GPT-DFSDT.
-For every LLM, we adopt a temperature setting of 0 to ensure reproducibility.
-It is worth noting that the responses from the baselines, ToolLLaMa, and GPT-DFSDT, are in natural language format, whereas the answers in SoAyBench are the responses returned by SoAPI.
-In order to ensure a fair comparison, we have designed a set of prompts to employ ChatGPT to extract precise answers from responses in a natural language format.
+We focus on the following Research Questions: (RQ1) Is the SoAy method more effective than the baseline method in the task of utilizing the AMiner API to assist academic information retrieval?
+(RQ2) Is the SoAy method more efficient than the baseline method?
+(RQ3) Can the performance of the SoAy method in practical applications satisfy users?
+We first conduct a detailed comparison of the effectiveness of SoAy and previous methods on the SoAyBench test set using our carefully designed metric, SoAyEval.
+Then, we deploy an application by SoAy and validate its effectiveness on collected user queries.
+6.1. Experiments on SoAyBench
+6.1.1. Experimental Settings.
+We propose a detailed and interpretable evaluation method based on SoAyBench and conduct comparative experiments with representative baselines.
 SoAyEval.
-Given a question that yields an API solution along with an invoked answer, we have outlined five types of evaluation metrics.
-These metrics take into account the accuracy of the solution, program, and answer together: (EM): Both the retrieved solution and answer Exactly Match the ground truth.
-(DS): The answer is correct, but a Different Solution is retrieved compared to the ground truth.
+Enabling large models to use tools for academic information retrieval is typically a multi-step task, with different steps often requiring the model to demonstrate different capabilities.
+We need detailed evaluation metrics that can help us compare the strengths and weaknesses of different methods and models in specific steps and capabilities.
+We propose SoAyEval, a detailed evaluation method designed for SoAyBench.
+Given a (Query, Solution, Code) triplet from SoAyBench, SoAyEval first executes the Code in AMiner Environment to get the ground truth answer, then compares the tested result with these ground truth solutions, code, and answer.
+(EM): Both the solution and answer Exactly Match the ground truth.
+(DS): The answer is correct, but a Different Solution is generated compared to the ground truth.
 (WS): The answer is wrong due to a Wrong Solution.
-(WP): The solution is correct but the answer is wrong, due to a Wrong Program generated for the solution, which can be executed but yields the wrong answer.
-(EE): Execution Error, which may by caused by the generation of a non-executable program or network errors during the APIs request.
+(WC): The solution is correct but the answer is wrong, due to a Wrong Code generated for the solution, which can be executed but still leads to a wrong answer.
+(EE): Execution Error, which may be caused by the generation of a non-executable code or network error during the request.
 Based on the aforementioned metrics, we define the accuracy metric as follows:
-[…]
+$\text{ACC}=\text{EM}+\text{DS},$
 which represents the proportion of correct answers produced by the method under test.
@@
 $\text{Score}=\frac{w_{1}\cdot\text{ACC}_{1}+w_{2}*\cdot\text{ACC}_{2}+w_{3}\cdot\text{ACC}_{3}}{w_{1}+w_{2}+w_{3}}.$
-We name this evaluation method SoAyEval.
-SoAyEval meticulously analyzes the entire process of large models using a set of APIs to answer corresponding questions.
-It is a universal evaluation method that could be adopted by other API-using scenarios.
 Notably, we assign $w_{1}=1,w_{2}=2,w_{3}=3$ in our experiment to highlight the multi-hop question-answering ability of the methods under test.
-7.2 Overall Performance (RQ1, 2)
-Table 3 presents the experimental results of SoAyGPT, SoAyLLaMA, and the other two representative baseline methods on SoAyBench.
-SoAyLLaMA achieves the highest overall score, with Code-13B Model of 91.24%, followed by the other SoAy method, SoAyGPT, which gets 86.57% with GPT-4 as the backbone.
-These results prove the effectiveness of SoAy.
-With the help of the prepared solution library, SoAy describes the complex inter-dependency and diverse execution logic to the LLMs.
-This significantly reduces the pressure of LLMs when determining the optimal execution path from the coupling APIs.
-Despite undergoing specialized tool usage training, ToolLLaMA does not score highly.
-This is primarily because ToolLLaMA’s training objective is to master the use of general APIs, where the dependencies involved are significantly less complex than those in complicated academic APIs.
-Furthermore, many questions in SoAyBench require a solution that incorporates branch or loop logic.
-Facing with the need to execute these complicated execution logics, GPT-DFSDT’s decision-tree-based inference encounters numerous repetitive sub-nodes, making the solution exceptionally challenging.
-7.3 Error Analysis (RQ1)
-The detailed analysis of model performances from SoAyEval reveals distinct characteristics and potential improvement areas for each model across different question types.
-The SoAyLLaMA and SoAyGPT models, particularly their latest versions, demonstrate superior capabilities in managing complex question types with minimal errors.
-ToolLLaMA and early versions of GPT-DFSDT show greater susceptibility to execution and solution pathway errors, indicating potential areas for algorithmic or training enhancements.
-Overall, advancements in model versions (notably from GPT-DFSDT 3.5 to 4 and SoAyGPT 3.5 to 4) show tangible improvements in reducing specific error types and enhancing overall accuracy, highlighting the benefits of continuous model development and refinement.
-7.4 Efficiency (RQ3)
-We evaluate the efficiency of the proposed method by reporting the average response time of various methods evaluated on SoAyBench, as shown in Table 4.
-We apply both SoAyLLaMA and ToolLLaMA in a network request service format to keep justice with GPT-based methods which should only be accessed through APIs calls.
-Both ToolLLaMa and GPT-DFSDT rely on multiple executions to complete complex API operations, thus consuming a considerable amount of time.
-Implementation through SoAy can effectively reduce the inference time of tool calls.
-Even with GPT-4, the average execution time of SoAy is nearly one-third of the execution time of GPT-DFSDT with GPT-4.
-Moreover, with a smaller parameter scale, the lightweight SoAyLLaMA models show even less execution time, which we consider already meet the requirement for being applied to an academic searching service.
-Table 4: Average response time of different methods (second).
+Baselines.
+We select GPT-DFSDT, utilizing the same three GPT versions as SoAyGPT, and ToolLLaMA-7b (Qin et al., 2023b) as baselines.
+GPT-DFSDT represents the current state-of-the-art method on multiple general domain tool-utilizing benchmarks(Qin et al., 2023b; Guo et al., 2024).
+ToolLLaMA-7b is an open-source model fine-tuned on a large corpus of tool-utilizing data, achieving comparable results to GPT-DFSDT.
+Since there are no benchmark datasets or methods for evaluating the task of academic information retrieval using APIs, we select the two methods mentioned above, which respectively represent the most effective and efficient methods in general domain tool-using tasks.
+We believe these methods can provide relatively strong comparisons for validating the effectiveness of SoAy in the domain of academic information seeking.
+It is worth noting that results from ToolLLaMA and GPT-DFSDT are in natural language format.
+To ensure a fair comparison using SoAyEval, we instruct ChatGPT to extract precise answers from their responses.
+Also, ToolLLaMA and GPT-DFSDT don’t have a module to generate a solution or plan before inferring, so we extract the solution from their path of API calls.
+For all LLMs, a temperature setting of 0 is used to ensure reproducibility.
+Table 4. Average response time of different methods (second).
 Method 7B 13B 3.5 3.5-16k 4
@@
 SoAyLLaMA-Code 1.12 1.35 / / /
-7.5 Backbones Comparison (RQ4)
-It’s worth considering which backbone model to choose to construct an academic searching system whether through SoAyGPT or SoAyLLaMA.
-Based on the results of SoAyBench, if we focus on the effectiveness of the results, the choice of SoAyGPT definitely falls on the latest version (GPT-4 when the paper is drafted).
-As for the choice of SoAyLLaMA, whose implementation is described in Appendix B, we can find a trend where the scores increase as the number of parameters grows.
-Simultaneously, given that Soay implements the planning and formatting processes defined in the problem section through code generation, CodeLlama, which excels in code generation, outperforms ChatLlama.
-It should be mentioned that the effectiveness shown on the results of SoAyBench should not be the sole criterion for judgment when considering practical applications.
-8 Related Works
-APIs, short for Application Programming Interfaces, are the software interfaces that provide services to other software components [^1] https://en.wikipedia.org/wiki/API#cite_note-1.
-It facilitates the exchange of data between various software applications, thereby enabling interoperability between them Ofoeda et al.
-(2019).
-For the application of LLMs using the API-using methodology, the current solutions and technical approaches can be divided into two major groups: A straightforward idea is to broaden the tool/API-using coverage for LLMs.
-Such efforts not only include basic web services, such as search engines Nakano et al.
-(2021); Liu et al.
-(2023); Zhang et al.
-(2023) and calculators Schick et al.
-(2023), but also include the intelligent multi-modal functions, such as vision understanding Wu et al.
-(2023) and image generation Lu et al.
-(2023).
-Others are devoted to the tool’s dependencies, programming logic and planning ability King et al.
-(2023); Shen et al.
-(2023); Paranjape et al.
-(2023).
-Except for the efforts that regard LLMs as database managers Qin et al.
-(2023a), there are also plenty of methods that modeling the tool’s graphical relations Zhang (2023), building intermediate instructions Burnett and McIntyre (1995) and accomplish complex tasks Song et al.
-(2023).
-Besides, applying LLMs to directly access the data from the Database through techniques like text-to-SQL Qin et al.
-(2022) proves ineffective.
-The limitation arises from the unavailability of access to the original data sources of related systems.
-9 Conclusion
-In this work, we propose an API-Using methodology, SoAy, to apply LLMs to academic information searching.
-The core technique of SoAy is the high-quality automatic built dataset, SoAyBench, providing 3960 academic triplets data.
-We align LLMs to generate codes on pre-constructed guidance, which could face the coupling and execution logic diverse challenges.
-An evaluation strategy, SoAyEval, is proposed to evaluate the AMiner API-using capability.
-Experiments show SoAy’s effectiveness, efficiency, and usability.
-In the future, we will continue to enlarge the SoAyBench and apply SoAy to more scenarios and domains.
-10 Limitations and Future Directions
-In this paper, we propose an approach that combines automated data generation with LLM alignment for academic query API-based tasks and demonstrates its effectiveness.
-The released SoAyBench and the proposed SoAyEval could support for evaluating related methods in subsequent studies.
-However, there are still some worthwhile explorations that we have not conducted:
-  • SoAyBench is a large-scale dataset constructed through automation, consisting of 3960 triplet data instances.
-Although we incorporate quality checks in the solution library construction, query generation, and code generation processes, and have human experts inspect and refine the test set, we must acknowledge that we cannot guarantee the provided tests in SoAyBench align perfectly with people’s expectations for the API system.
-Some prior worksQin et al.
-(2023b) validate the quality of their benchmarks by comparing the results of tested methods on their own benchmark with manually evaluated results.
-This is an aspect we plan to further explore and validate in the future, which may lead to more human-aligned automated constructed datasets.
-  • The API-using approach we propose addresses the task of guiding large models to utilize complex academic search APIs.
-However, we have not yet validated whether our method has advantages in simpler scenarios.
-We will continue to expand SoAyBench and conduct practicality and effectiveness validations in various scenarios.
-  • In practical applications, we have encountered an inconvenient point of updating the deployed system: if we want to incorporate new APIs on top of existing ones, especially when adding a large number of APIs, the context-guided SoAyGPT needs to adjust the structure of the prompt within a limited input space.
-This poses a challenge that requires a trade-off between the breadth and granularity of examples.
-Although our proposed training-based SoAyLLaMA theoretically allows for the unlimited expansion of APIs by adding sufficient training data, this is an aspect we have not explored yet.
-In the future, we plan to explore methods for continuous learning with SoAyLLaMA to enable rapid inclusion of new API data without compromising the model’s mastery of existing APIs.
-In conclusion, the quality of data generation, scalability, and the scenarios presented in this paper opens up numerous possibilities for further exploration.
-In the future, we will continue to fulfill the current work and explore further in these mentioned directions.
-11 Acknowledgements
-This project would like to thank the data annotators at Zhipu AI for their help and support. All GPU compute and API costs are provided by Zhipu AI.
-12 Ethical Considerations
-12.1 New datasets
-Using SoAy, our new dataset is built on the AMiner dataset.
-The construction of this new dataset relies on the efforts of annotators.
-We value their contributions and provide them with reasonable compensation based on the quantity of annotations and the duration of their work.
-For the characteristics of the dataset, it focuses on the academic field and is suitable for academic researchers.
-Furthermore, as talked before, to prevent additional risks, we establish an evaluation benchmark, which is further enhanced by additional human verification to ensure the quality of this new dataset and its application effectiveness.
-12.2 NLP application
-The intended use of the SoAy trained LLMs is to facilitate researchers in better accessing academic literature and other related resources.
-It allows for querying in natural language, reducing the technical barriers for scholars and others with academic information retrieval needs.
-However, just like the issues inherent to LLMs, sometimes it may provide incorrect information, potentially leading to misguidance for users, which may be a serious disruption for researchers since academic information needs to be accurate.
-Appendix
-In the appendix, we’ll present in three folds the details of SoAy and SoAyBench, along with more experimental details and conclusions.
-Part I focus on the implementation of SoAy, including the details of SoAyGPT in  A, SoAyLLaMA in  B and SoAyBench in  C.
-Part II shows more about the experiments, including the online evaluation in  4, the implementation of SoAy on other academic information seeking systems in  F, and more interesting findings from the results on SoAyBench in  E.
-Part III lists all the prompts used in SoAy and SoAyGPT.
-Part I. Implementation Details
-The content of this section primarily introduces some implementation details of the SoAy method, including the inference architecture of SoAyGPT, the training details of SoAyLLaMA, as well as the data details and API environment setup details of SoAyBench.
-We implement SoAyGPT on three distinct versions of GPT, specifically 3.5: gpt-3.5-turbo-0613; 3.5-16k: gpt-3.5-turbo-16k-0613, and 4: gpt-4-0613.
-We conduct the fine-tuning on 3 different variants of the LLaMA model, respectively Chat-7B: Llama-2-7b-chat-hf, Code-7B: CodeLlama-7b-Instruct-hf, and Code-13B: CodeLlama-13b-Instruct-hf.
-Appendix A SoAyGPT
-SoAyGPT is an alignment method designed for a series of large-scale models that are closed-source yet possess high command-following capabilities.
-Specifically, SoAyGPT employs in-context learning to construct three LLM agents Wang et al.
-(2024) (Solution Retrieval Agent, Code Generation Agent, and Answer Generation Agent) and an environment operable by these agents to complete the alignment process.
-The agents and environments are designed:
-Solution Retrieval Agent.
-Given the user input $Q$, the solution retrieval agent obtains the appropriate solution from the solution library $S$.
-It’s important to note that the retrieval process here diverges from traditional information retrieval functions, which extract the exact required element from a set.
-Essentially, this retrieval module presents all possible solutions as the prompt to the LLMs, enabling them to generate a suitable solution.
-Consequently, it’s possible that the produced solution differs from any solution in $L_{S}$.
-Code Generation Agent.
-The code generation agent must generate the correct program under the guidance of the retrieved solution $S$.
-Since the solution might not be included in $L_{S}$, if the retrieved solution $S$ is in $L_{S}$, we simply use $P_{p}$ specifically defined for $S$ to generate the required program.
-Otherwise, SoAy has to use $P_{g}$ that is defined for all the possible solutions to guide the LLMs in generating a program, as no specific solution could be identified.
-SoAy employs Python as the programming language for each program, allowing it to incorporate not only sequential but also branching and looping logic.
-Code Execution Environment.
-SoAy employs a virtual Python environment as the executor to obtain the results, serving as the knowledge $K$.
-Here, $K$ is expected to be the exact answer to $Q$ if $Q$ asks about some precise information, like the number of citations for a scholar.
-Otherwise, $K$ provides helpful information if $Q$ has an ambiguous intent.
-Answer Generation Agent. Finally, SoAy generates a natural language answer $A$ based on $Q$ and $K$, guided by $P_{a}$.
-Appendix B SoAyLLaMA
-SoAyLLaMA is designed for supervised fine-tuning of open-source models when training resources are available.
-Through the training stage, SoAyLLaMA integrates an API description and queries as inputs and combines solutions and programs as responses.
-During the inference stage, SoAyLLaMA concurrently completes the Planning and Formatting processes to derive solutions and code.
-Similar to SoAyGPT, it utilizes the environment to execute and obtain answers based on $K$.
-We conduct the fine-tuning on 3 different variants of LLaMA Model, respectively Chat-7B: Llama-2-7b-chat-hf [^2] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, Code-7B: CodeLlama-7b-Instruct-hf [^3] https://www.modelscope.cn/models/AI-ModelScope/CodeLlama-7b-hf/summary, Code-13B: CodeLlama-13b-Instruct-hf [^4] https://huggingface.co/meta-llama/CodeLlama-13b-Instruct-hf.
-Dataset. As mentioned in Section 3, four-fifth of the instances from SoAyBench are utilized here to train the SoAyLLaMA. We simply take the academic query as an input and the relevant solution.
-Training Details.
-We train 13B models on 8 * NVIDIA A100 GPU (80GB) for 13 hours to accomplish a convergent loss, with a batch size of 8.
-7B models are trained on 4 * NVIDIA A100 GPU (80GB) for 7 hour with a batch size of 16.
-All the training details are saved in the training log in our Github Repo [^5] https://anonymous.4open.science/r/SoAy-C025/readme.md.
-Appendix C SoAyBench
-To assess AMiner API-using capabilities, the foundational AMiner APIs for LLMs to invoke and a test set should be provided.
-However, given the dynamic and specialization nature of academic data, with academic entities-related information rapidly changing, maintaining a test set with static answers proves challenging.
-To address this challenge, we clone AMiner’s APIs at a specific point in time to create a static service, from which we create a corresponding static test set including 792 meaningful academic questions (one-fifth of the whole SoAyBench) and ground truth conducted from the cloned APIs.
-We extract a subset of the comprehensive AMiner database, which includes over 66,000 entities such as persons and publications, at a specific point in time, September 23, 2023.
-This subset serves as our experimental evaluation environment, and we clone an API Service based on it.
-Consequently, the results returned by these APIs, given the same parameters, will remain consistent over time.
-This dataset and the API service provide a foundation for our subsequent construction of the test set.
-Part II. Experiments Details
-In this section, we will present more details related to the experiments.
-To verify the efficiency and effectiveness of SoAy in real applications, it is insufficient to rely solely on its performance on SoAyBench, so we designed an online evaluation to test SoAy’s performance in real-use scenarios.
-Beyond the research questions in Section 7, more interesting insightful conclusions can be found in the results on SoAyBench.
-Furthermore, to validate the universality of the SoAy method, rather than it only working for the AMiner system, we implemented SoAyGPT in two other academic information seeking scenarios and demonstrated its results in a case study manner.
-Appendix D Online Evaluation
-Figure 4: Distribution of questions used for online human evaluation.
+6.1.2. Overall Performance (RQ1).
+Table 3 presents the experimental results of SoAyGPT, SoAyLLaMA, and the other two representative baseline methods on SoAyBench.
+SoAyLLaMA achieves the highest overall score, with Code-13B Model of 92.74%, followed by the other SoAy method, SoAyGPT, which gets 86.57% with GPT-4 as the backbone.
+These results prove the effectiveness of SoAy.
+With the help of the prepared solution library, SoAy describes the complex inter-dependency and diverse execution logic of the LLMs.
+This significantly reduces the pressure of LLMs when determining the optimal execution path from the coupling APIs.
+Despite undergoing specialized tool usage training, ToolLLaMA does not score highly.
+This is primarily because ToolLLaMA’s training objective is to master the use of general APIs, where the dependencies involved are significantly less complex than those in complicated academic APIs.
+Furthermore, many questions in SoAyBench require a solution that incorporates branch or loop logic.
+Facing the need to execute these complicated execution logic, GPT-DFSDT’s decision-tree-based inference encounters numerous repetitive sub-nodes, making the solution exceptionally challenging.
+6.1.3. Efficiency (RQ2).
+We evaluate the efficiency of the proposed method by reporting the average response time of various methods evaluated on SoAyBench, as shown in Table 4.
+We apply both SoAyLLaMA and ToolLLaMA in a network request service format to keep justice with GPT-based methods which should only be accessed through API calls.
+Both ToolLLaMA and GPT-DFSDT rely on multiple executions to complete complex API operations, thus consuming a considerable amount of time.
+Implementation through SoAy can effectively reduce the inference time of tool calls.
+Even with GPT-4, the average execution time of SoAy is nearly one-third of the execution time of GPT-DFSDT with GPT-4.
+Moreover, with a smaller parameter scale, the lightweight SoAyLLaMA models show even less execution time, which we consider already meets the requirement for being applied to a real-world service.
+Figure 4.
+Distribution of questions used for online human evaluation.
 Questions are categorized by 1, 2, and 3-hop types, further divided into exact and ambiguous categories.
-The ambiguous category is specific to this online evaluation setting.
-We examine the Spearman correlation coefficients among the voting results of the 15 human annotators, as shown in Figure 5.
-The released SoAyBench contains the templated questions derived from the AMiner SoAPI, thus testing a method’s ability to understand and utilize the SoAPIs to a certain extent.
-However, in real-world deployment scenarios, the goal of a good method should be to fulfill user needs more comprehensively and offer a superior user experience.
-To achieve this, we implement the proposed SoAy as an online application, gather 52 real user demands from the logs, and invite 10 annotators to conduct human evaluation.
-Figure 5: Spearman correlation among 15 annotators
-Settings.
-We deploy SoAy with GPT-3.5-16k as its backbone online, due to the superior speed of the 3.5 version, enhancing user experience.
-Over time, we gather 52 authentic questions from the application logs.
-The distribution of these questions is illustrated in Figure 4, categorized not only by their 1, 2, and 3-hop nature but also classified as either exact or ambiguous.
-Ambiguous questions, such as requests for “comparisons between two scholars”, lack standardized answers and thus yield diverse responses.
-We find that the majority of these 52 questions could be resolved within two hops.
-To evaluate the answers, we invite five academic experts to provide responses based on their expertise, supplemented by the AMiner system and general search engines.
-The online system’s answers, responses directly from ChatGPT, and the aforementioned human-provided answers constitute the comparison methods.
-For a balanced and fair assessment of the experimental results, we develop an annotation platform and create a detailed annotation manual.
-This platform allows annotators to shuffle and anonymize the answers from the three methods before each voting session.
-A team of 10 annotators is recruited to assess the answers provided by each method for every question across a two-day period.
-Their focus is on evaluating the accuracy and readability of the responses, culminating in a vote for the most effective response among the three.
-We also open-source the platform and its associated manual.
-Evaluation Results.
-Figure 6: Voting results for authentic online questions.
-After identifying the three annotators with the lowest correlation to others, we exclude their input.
-For the remaining annotators, a majority voting strategy is employed to consolidate the votes for the three methods.
-The summarized results are presented in Figure 6.
-These results show a preference for the SoAy method’s responses in real-world scenarios across various question types, outperforming both direct ChatGPT and human expert responses.
-Notably, SoAy’s responses are particularly favored for questions with exact answers.
-However, SoAy’s advantage is less pronounced for ambiguous questions when compared to baseline responses.
-This observation aligns with SoAy’s design and the assessment objectives of SoAyBench, both of which prioritize precise answering through the integration of SoAPI Library resources.
-Nevertheless, there’s a noticeable gap between the nature of real-world inquiries and the templated questions in SoAyBench, posing a challenge for SoAy in addressing certain scenarios.
-This finding points towards the potential for integrating SoAPI calls with model-intrinsic functions for a more robust response system.
-Appendix E Insights on SoAyBench
-E.1 Intriguing phenomena
+The ambiguous category is specific to the online evaluation setting.
+[…]
+Distribution of questions used for online human evaluation.
+Questions are categorized by 1, 2, and 3-hop types, further divided into exact and ambiguous categories.
+The ambiguous category is specific to the online evaluation setting.
+6.1.4. Variances.
+Notably, the data fluctuations shown in Table 3 do not reflect the variations that would result from repeating the same experiment on identical samples.
+Given that the $768$ test samples in SoAyBench originate from $44$ template questions, and each template is filled with different entities to create unique question instances, we compute each metric within a group of instantiated questions from the $44$ question templates.
+We then calculate the mean and variance across groups of different instantiated questions.
+Our observations indicate that SoAyGPT-4 and SoAyLLaMA deliver a more consistently stable performance across various error types, which leverages GPT-4’s improved generalization ability and the high-quality training data provided by SoAyBench.
+6.1.5. Intriguing phenomena
 Table 3 reveals several other intriguing phenomena: First, while SoAy excels with GPT-4, its advantages are less marked on GPT-3.5-turbo and GPT-3.5-16k (hereafter collectively referred to as GPT-3.5 backbones), trailing behind the GPT-DFSDT method with the same GPT-3.5 backbones.
-Notably, the DS and WS metrics for SoAy with GPT-3.5 backbones are quite high.
+Notably, the DS and WS metrics for SoAy with GPT-3.5 backbones are quite high (the lower, the better).
 This can be attributed to GPT-4’s superior instruction-following proficiency, which amplifies the impact of SoAy’s complex prompts on the model adept at instruction following.
@@
 Third, GPT-DFSDT exhibits a stronger understanding of one-hop questions when using GPT-3.5 backbones.
-It breaks down complex questions into simpler one-hop questions, determining the appropriate APIs and their parameters step by step, which enhances its comprehension of one-hop questions.
-E.2 Variances
-Notably, the data fluctuations marked in Table 3 do not represent variances represented by repeating the identical experiment on identical samples.
-Given that the $768$ test samples in SoAyBench originate from $44$ template questions, and each template is filled with different entities to create unique question instances, we compute each metric within a group of instantiated questions from the $44$ question templates.
-We then calculate the mean and variance across groups of different instantiated questions.
-Our observations indicate that SoAyGPT-4 and SoAyLLaMA delivers a more consistently stable performance across various error types, which benefits from GPT-4’s enhanced generalization capacity across a variety of entities and sufficient and diverse training data from SoAyBench.
-E.3 Detailed Error Analysis
-Notably, We do not rectify network errors because SoAy aims to apply the system in real-world use, implying that it is essential to test each model’s stability.
-During experiments, we attempt to request the models up to 3 times.
-If all 3 attempts fail, the situation is classified as an EE.
-E.3.1 ToolLLaMA (7B)
-Shows the highest Execution Error (EE) rates across all question types, which might indicate issues with the model’s ability to generate executable programs consistently.
-It struggles particularly with one-hop questions, which have the lowest Exact Match (EM) rate and high Wrong Solution (WS) errors.
-Two-hop and three-hop questions reflect similar challenges with WS and EE, although DS errors decrease slightly as the complexity increases.
-High variability in error rates suggests that the model might benefit from training focused on error reduction techniques.
-The overall score of 16.72 is the lowest among the models, reflecting significant room for improvement.
-E.3.2 GPT-DFSDT
-Version 3.5: One-hop questions see extremely high DS rates but lower WS and EE, suggesting issues with solution path diversity but not necessarily with execution.
-Two-hop and three-hop configurations show a balance of errors across DS, WS, and EE, with a slightly better EM rate for three-hop.
-Version 3.5-16K: This model shows improved performance over the standard 3.5 version, with particularly notable improvements in EM and ACC for one-hop questions.
-DS errors are significantly reduced compared to the 3.5 version, indicating enhanced solution alignment.
-Version 4: The most advanced version, showing the best performance with notably high scores, especially in one-hop questions which almost reach perfect EM and ACC.
-Lower WS rates across all configurations, indicating robustness in generating correct solutions.
-E.3.3 SoAyGPT
-Version 3.5: Generally strong in EM and ACC, particularly for one-hop questions, suggesting effective solution generation and execution.
-Lower EE than ToolLLaMA, indicating more reliable program execution.
-Version 3.5-16K: Consistently high EM rates across question types, with particularly low DS errors in two-hop questions. Shows good balance across all error types, with no extreme weaknesses.
-Version 4: Outstanding performance in one-hop questions with nearly no errors and very high EM and ACC.
-Two-hop and three-hop questions also show very high scores, making this version highly effective and reliable.
-E.3.4 SoAyLLaMA
-Chat-7B: Exhibits nearly perfect performance in one-hop questions with zero DS, WS, and WP errors.
-Three-hop questions maintain high performance but see a slight increase in WP errors, which still remain low compared to other models.
-Code-7B: Solid performance with minimal error rates across all question types. A slight uptick in WP errors in two-hop questions could be a focal point for model refinement.
-Code-13B: The best-performing model across all metrics, with impressive EM and ACC in one-hop questions and consistent strength in multi-hop configurations.
-Nearly negligible error rates, setting a high standard for large language models.
-Appendix F Other Scenarios
-F.1 SoAy implemented on CrossRef
-F.1.1 Introduction
-Crossref is a registration agency for DOIs (Digital Object Identifiers) established by a consortium of renowned global publishers.
-It registers DOI numbers for newly published and previously published journal articles.
-All works registered with Crossref by various publishers automatically establish citation links with each other, facilitating readers to navigate journals and access the websites of different publishers, thus significantly increasing the chances of journal articles being accessed and cited.
-Crossref’s open reference linking system establishes citation links to journal articles between different publisher websites.
-For developers, Crossref provides rich open-source API interfaces on GitHub [^6] https://github.com/CrossRef/rest-api-doc?tab=readme-ov-file, such as search and other functionalities.
-F.1.2 Implementation
-The codes, prompt, and APIs of this part of the implementation are available at our GitHub repo: https://github.com/RUCKBReasoning/SoAy/tree/master/scenarios/CrossRef.
-As shown above, we need to finish two primary stages: the offline solution library preparation stage and the online solution-based inference stage.
-Figure 7: Illustration of Crossref SoAPI Library, SoAPI Graph and solution library.
-API Library.
-As shown in Figure 7, we use 6 APIs of Crossref.
-The phenomenon of API coupling is prevalent among these SoAPIs.
-For example, the parameter doi for getWorksByDoi(·) must first be obtained by searchWorkByTItle(·).
-This necessitates a specific execution order where searchWorkByTItle(·) must be used before getWorksByDoi(·), and a specific execution composition where if getWorksByDoi(·) is used, searchWorkByTItle(·) is definitely used.
-Solution Library Construction.
-The solution library is the cornerstone of the entire framework.
-Just as what we did for AMiner, we finish the solution library for Crossref, proving SoAy’s usability.
-The initial stage of SoAy involves analyzing Crossref’s SoAPI and constructing the SoAPI Graph.
-Following this, a process of sampling and filtering paths is applied to the SoAPI Graph, resulting in the solution library.
-Based on it, the prompt library is then constructed.
-To better understand the performance of Soay on Crossref, let’s take a detailed look at Crossref’s prompts and cases on Figure 10
-General Prompt Drafting.
-The General Prompt outlines a structured approach to crafting programs that utilize a set of APIs to address various queries related to literature and authors.
-It starts by presenting the APIs that will be employed, such as ‘searchBook‘, ‘getBook‘, and others, which are instrumental in searching for books, obtaining book details, and accessing author information.
-The prompt then proceeds to illustrate the process through a series of examples.
-The first example guides the user on how to use the ‘searchBook‘ API to find publishers in a particular subject area (XXX), iterating through the results to compile a list of publisher names.
-This demonstrates the initial step of narrowing down the scope of the query to a more manageable list.
-Following that, the second example builds upon the first by not only identifying publishers in a specific field but also by delving deeper to list articles or works associated with those publishers.
-This is achieved by chaining the ‘searchPublisherBySubject‘ API to ‘getPublisherBasicInfo‘ and then to ‘getPublisherWorks‘, showcasing a more complex use of the APIs to extract detailed information.
-The prompt concludes with a user query placeholder, encouraging the LLMs to apply the learned methods to formulate a solution.
-This involves selecting the correct sequence of APIs to retrieve the necessary information in response to the query.
-The prompt is designed to be a comprehensive guide that adapts to different levels of complexity in the user’s requests, from simple to more advanced searches.
-F.1.3 Case Study
-Case 1: Mathematics Publishers.
-The first case presents a query for a list of publishers in the mathematics field.
-The solution uses the ‘searchPublisherBySubject‘ API with ’math’ as the input to extract a list of publishers, along with the number of DOIs they’ve published.
-The result provides a collection of publisher names and their respective DOI counts, highlighting their impact in the field of math.
-Case 2: DOI Numbers in Computer Science.
-The second case addresses the request for current DOI numbers from publishers in computer science.
-By utilizing ‘searchPublisherBySubject‘ to identify relevant publishers, followed by ‘getPublisherBasicInfo‘ to obtain DOI counts, the solution compiles a list of publishers with their current DOI numbers, showcasing their recent publication activities.
-Case 3: Article Titles in Computer Science.
-The third case involves listing articles in the computer science domain.
-It starts with ‘searchPublisherBySubject‘ to find publishers, then uses ‘getPublisherBasicInfo‘ and ‘getPublisherWorks‘ to retrieve a list of article titles.
-The output is a concise list of article titles that represent the range of research within computer science.
-F.2 SoAy implemented on OpenLibrary
-F.2.1 Introduction
-Open Library is an online project aimed at creating a comprehensive database of all published books.
-Operated by the Internet Archive, the website has been serving the public since 2006.
-It provides information pages for thousands of books, displaying their summaries, covers, publication details, and more.
-Additionally, it features personal pages for numerous authors, detailing their biographical information, major works, and areas of expertise.
-For developers, the website also offers web APIs such as search functionality, which are documented in the Public APIs project.
-F.2.2 Implemention
-The codes, prompt, and APIs of this part of the implementation are also available at our GitHub repo: https://github.com/RUCKBReasoning/SoAy/tree/master/scenarios/OpenLibrary.
-The same as AMiner and CrossRef, we need to finish two primary stages: the offline solution library preparation stage and the online solution-based inference stage.
-Figure 8: Illustration of OpenLibrary SoAPI Library, SoAPI Graph and solution library.
-API Library.
-As shown in Figure 8, we use 6 APIs of Crossref.
-The phenomenon of API coupling is prevalent among these SoAPIs.
-For example, the parameter subject for searchSubject(·) must first be obtained by getBook(·) or getAuthorBasicInfo(·).
-This necessitates a specific execution order where getBook(·) or getAuthorBasicInfo(·) must be used before searchSubject(·), and a specific execution composition where if searchSubject(·) is used, getBook(·) or getAuthorBasicInfo(·) is definitely used.
-Solution Library.
-The solution library is the cornerstone of the entire framework.
-Just as what we did for AMiner, we finish the solution library for OpenLibrary, proving SoAy’s usability.
-The initial stage of SoAy involves analyzing OpenLibrary’s SoAPI and constructing the SoAPI Graph.
-Following this, a process of sampling and filtering paths is applied to the SoAPI Graph, resulting in the solution library.
-Based on it, the prompt library is then constructed.
-General Prompt.
-To better understand the performance of Soay on OpenLibrary, let’s take a detailed look at OpenLibrary’s prompts and cases on Figure 11 It starts by presenting the APIs that will be employed, such as ‘searchBook‘, ‘getBook‘, and others, which are instrumental in searching for books, obtaining book details, and accessing author information.
-The prompt then proceeds to illustrate the process through a series of examples.
-The first example guides the user on how to use the ‘searchBook‘ API to find publishers in a particular subject area (XXX), iterating through the results to compile a list of publisher names.
-This demonstrates the initial step of narrowing down the scope of the query to a more manageable list.
-The same principle applies to the two other examples.
-The prompt concludes with a user query placeholder, encouraging the LLMs to apply the learned methods to formulate a solution.
-This involves selecting the correct sequence of APIs to retrieve the necessary information in response to the query.
-The prompt is designed to be a comprehensive guide that adapts to different levels of complexity in the user’s requests, from simple to more advanced searches.
-F.2.3 Case Study
-Case 1: Identifying Mathematics Publishers.
-In the first case, the objective is to identify publishers that specialize in mathematics.
-The process begins by utilizing the ‘searchPublisherBySubject‘ API, specifying ’math’ as the subject parameter.
-This call returns a list of publishers active in the field of mathematics, each accompanied by the number of DOIs (Digital Object Identifiers) they have generated.
-The output is a curated list of publisher names alongside their DOI counts, which serves as an indicator of their contributions and presence in the mathematical literature space.
-Case 2: Obtaining DOI Numbers for Computer Science Publishers.
-The second case focuses on obtaining the current DOI numbers for publishers within the computer science field.
-The ‘searchPublisherBySubject‘ API is employed to first filter out publishers related to computer science.
-Once a list of these publishers is obtained, the ‘getPublisherBasicInfo‘ API is then used to fetch the up-to-date DOI numbers for each publisher.
-The result is a compilation of publisher names paired with their current DOI counts, providing a snapshot of their recent scholarly output and engagement within the field.
-Case 3: Listing Computer Science Articles.
-The third case aims to extract a list of article titles from the computer science domain.
-The process initiates with the ‘searchPublisherBySubject‘ API to pinpoint publishers in the field, followed by the use of ‘getPublisherBasicInfo‘ to acquire the publisher’s unique ID.
-With the publisher ID in hand, the ‘getPublisherWorks‘ API is then applied to retrieve a comprehensive list of article titles associated with the publisher.
-The final output is a succinct array of article titles, offering a glimpse into the diverse range of research topics and findings published within the realm of computer science.
-Part III. Prompts
-In this section, we will comprehensively present the prompts used in the data generation process of SoAy, as well as the implementation process of SoAyGPT on AMiner and two other scenarios.
-Figure 9: A Preview of the prompt library.
-$P_{s}$ signifies the prompt for solution retrieval, $P_{p}$ represents the prompt for program generation, and $P_{a}$ denotes the prompt for answer generation.
-All the black text except the ellipsis is the original prompt content, the yellow highlighted text is the annotations added for better understanding in the context of creating a prompt, and the purple highlighted text is the content waiting for the LLM output.
-Figure 10: A Preview of the prompt library and three cases for Crossref.
-Figure 11: A Preview of the prompt library and three cases for OpenLibrary.
-Figure 12: A Preview of the prompts used in the data preparation process.
-Figure 13: The screenshots of the demonstration systems https://soay.aminer.cn and online service widget https://chatglm.cn/main/gdetail/65bf5a99396389a73ace9352 powered by SoAy.
+This is because the general idea of GPT-DFSDT involves breaking down complex questions into simpler one-hop questions, determining the appropriate APIs and their parameters step by step, which enhances its comprehension of one-hop questions.
+6.2. Deployment Evaluation
+Experiments on SoAyBench validate the effectiveness and efficiency of SoAy.
+However, in practical applications, the questions that users are truly concerned with may differ from those in the dataset.
+To test the user experience of SoAy in practical applications, we conduct the following deployment experiment.
+6.2.1. Settings.
+We deploy SoAy as a web application, then gather 52 representative real user queries from the logs, and invite 15 annotators to conduct human evaluation.
+Queries Detailes.
+The distribution of the 52 gathered queries is illustrated in Figure 4.
+We classify them not only into 1, 2, and 3-hop types as SoAyBench does but also into either exact or ambiguous questions.
+Ambiguous questions, such as requests for “comparisons between two scholars”, lack standardized answers and thus yield diverse responses.
+Figure 5. Voting results for authentic online questions.
+[…]
+Voting results for authentic online questions.
+Baselines.
+We compare the responses from SoAy, GPT-4, and human expert annotated answers on these 52 questions.
+Here GPT-4 indicates GPT-4-1106-preview and SoAy indicates SoAyGPT with GPT-4 as the backbone.
+We do not include GPT-DFSDT, ToolLLaMA, or other tool-using works as comparisons because, based on the results from Section 6.1.3, their efficiency clearly cannot support the requirements of practical deployment.
+Metrics.
+We invite 15 annotators to choose their favorite answer from three for each query.
+For a balanced and fair assessment of the experimental results, we develop an annotation platform and create a detailed annotation manual.
+This platform allows annotators to shuffle and anonymize the answers from the three methods before each voting session.
+This task did not require the annotators to have expert-level knowledge, so we recruit annotators who are interested in academic tasks through public channels.
+As such, the annotators’ level of diligence and cognitive abilities might vary significantly, and we aim to filter out annotators who did not annotate seriously, or those whose opinions deviate significantly from the mainstream, through consistency analysis.
+We examine the Spearman correlation coefficients among the voting results of the 15 human annotators, as shown in Figure 6.
+We sum up all the coefficients of each annotator to obtain a trustworthiness score for this annotator.
+We then select the top 10 annotators based on this score and use their votes for further analysis.
+For each question, the method selected by the most participants (out of ten) receives one vote.
+The final vote totals for each method are calculated by summing the votes across all 52 questions.
+Figure 6. Spearman correlation among 15 annotators
+6.2.2. Results (RQ3).
+The summarized results are presented in Figure 5.
+These results show a preference for the SoAy method’s responses in real-world scenarios across various question types, outperforming both direct GPT-4 and human expert responses.
+Notably, SoAy’s responses are particularly favored for questions with exact answers.
+However, SoAy’s advantage is less pronounced for ambiguous questions when compared to baseline responses.
+This observation aligns with SoAy’s design and the assessment objectives of SoAyBench, both of which prioritize precise answering through the integration of AMiner API Library resources.
+Nevertheless, there’s a noticeable gap between the nature of real-world inquiries and the templated questions in SoAyBench, posing a challenge for SoAy in addressing certain scenarios.
+This finding points towards the potential for integrating API calls with model-intrinsic functions for a more robust response system.
+7. Related Works
+In the era of Large Language Models (LLMs) (Zhao et al., 2023), the task of academic information retrieval can be accomplished by teaching LLMs to use the APIs of established academic systems (dblp team, 2005; Giles et al., 1998; Noruzi, 2005; Tang et al., 2008; Sinha et al., 2015).
+Teaching LLMs to handle these APIs is not easy as LLMs’ abilities are limited by the extent of knowledge acquired during pre-training (Radford et al., 2019).
+For the LLMs API-using methodology (Qin et al., 2023a), the current technical approaches can be divided into two major directions:
+A straightforward idea is to broaden the tool or API-using coverage for LLMs.
+Such efforts have tried to teach LLMs to handle more and more colorful tools/APIs, from the most basic web services (Nakano et al., 2021; Liu et al., 2023b; Zhang et al., 2023; Lazaridou et al., 2022a; Komeili et al., 2021; Lazaridou et al., 2022b) to calculators (Schick et al., 2024; Das et al., 2024), KB accessing tools (Thoppilan et al., 2022; Patil et al., 2023; Hao et al., 2024; Zhuang et al., 2024), code interpreter (Lu et al., 2023; Gao et al., 2023; Pan et al., 2023; Wang et al., 2023), multi-modal functions (Wu et al., 2023; Sur´ıs et al., 2023; Yang et al., 2023a, b; Gao et al., 2024; Zhao et al., 2024b) and other scenarios (Zhao et al., 2024a; Yao et al., 2022; Inaba et al., 2023).
+Some other works have focused on the need to connect multiple tools, requiring LLMs to analyze and plan these relationships (King et al., 2023; Shen et al., 2023; Paranjape et al., 2023).
+Examples include Chameleon (Lu et al., 2023), which firstly retrieves relevant APIs based on the given query and then separately executes them to get the responses, ToolChain (Zhuang et al., 2023), which uses a decision tree for planning, RestGPT (Song et al., 2023), which combines coarse and fine-grained planning, and ControlLLM (Liu et al., 2023a), which reasons based on a graph structure.
+These works often use prompt-based methods with limited manual examples, restricting LLMs’ reasoning potential, and can be inefficient due to the high number of reasoning steps needed.
+Training-based methods like ToolLLM (Qin et al., 2023b), ToolLink (Qian et al., 2023), TPTU-v2 (Kong et al., 2023), and OpenAGI (Ge et al., 2023) can teach LLMs to fully understand some APIs as there’s no limitation in the scale of training examples.
+However, these works focus more on APIs that can be individually used like weather query APIs and calculator APIs, but ignore the coupling between the APIs.
+Our method, SoAy, uses both tuning-free and training-based approaches to solve the coupling problem in academic APIs to seek academic information.
+8. Conclusion
+In this work, we propose an API-using methodology, SoAy, to apply LLMs to academic information searching.
+SoAy generates a solution firstly, which is a sequence of APIs, and then uses this solution to guide the code generation process.
+By parsing the solution, SoAy can handle the coupling relation between a set of APIs.
+Code generation and execution process makes SoAy more efficient than previous methods using multi-step reasoning.
+The core technique of SoAy is the high-quality automatic built dataset, SoAyBench, providing 3,960 academic triplets data.
+Based on the training set of SoAyBench, we build SoAyLLaMA and SoAyGPT, separately using the tuning method and tuning-free method.
+To evaluate the AMiner API-using capability of different models and methods, we propose a comprehensive evaluation strategy, SoAyEval.
+Experiments show SoAy’s effectiveness, efficiency, and usability.
+SoAy has been applied to real-world service and has been accessed more than 54,800 times.
+The implementation of SoAy to various applications in academic domain highlights the usability of SoAy.
+In the future, we will continue to enlarge the SoAyBench and apply SoAy to more scenarios and domains.
+In the future, we will continue to enlarge the SoAyBench and enrich the queries in SoAyBench based on the real-world application user logs.
+Acknowledgements.
+This work is supported by the National Key Research & Develop Plan (2023YFF0725100) and National Natural Science Foundation of China (62322214, U23A20299, U24B20144, 62172424, 62276270).
+This project would like to thank the data annotators at Zhipu AI for their help and support.
+All GPU compute and API costs are provided by Zhipu AI.
+Appendix
+Appendix A Application of SoAy
+As of the submission date, SoAy has been deployed for 10 months and has been accessed more than 54,800 times in the real-world application at https://www.aminer.cn/ and https://chatglm.cn/main/gdetail/65bf5a99396389a73ace9352, providing users with accurate academic information and helping them stay informed with up-to-date insights throughout their research process.
+Appendix B SoAyGPT
+SoAyGPT is an alignment method designed for a series of large-scale models that are closed-source yet possess high instruction following capabilities.
+We implement SoAyGPT on three distinct versions of GPT, in particular 3.5: gpt-3.5-turbo-0613, 3.5-16k: gpt-3.5-turbo-16k-0613, and 4: gpt-4-0613.
+Specifically, SoAyGPT employs in-context learning to construct three LLM agents (Wang et al., 2024) (Solution Generation Agent, Code Generation Agent, and Answer Generation Agent) and an environment operable by these agents to complete the alignment process.
+Identifying the solution first and then generating the code is less challenging for LLMs, compared to directly code generation, reducing the likelihood of producing incorrect answers.
+Additionally, the advantage of generating the solution and code separately, rather than simultaneously, is that each step allows for more information and examples to be provided within a limited input space.
+The agents and environments are designed as follows. Prompts for each agent are shown in Figure 7.
+Appendix C SoAyLLaMA
+SoAyLLaMA is designed for supervised fine-tuning of open-source models when training resources are available.
+During the training stage, SoAyLLaMA uses the API description and the query as input for each data sample, and combines the solution and code as the response.
+During the inference stage, SoAyLLaMA concurrently completes the Planning and Formatting processes to derive solutions and code.
+Similar to SoAyGPT, it utilizes the environment to execute and obtain answers based on $K$.
+We conduct the fine-tuning on 3 different variants of LLaMA Model, respectively Chat-7B: Llama-2-7b-chat-hf [^1] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf, Code-7B: CodeLlama-7b-Instruct-hf [^2] https://www.modelscope.cn/models/AI-ModelScope/CodeLlama-7b-hf/summary, Code-13B: CodeLlama-13b-Instruct-hf [^3] https://huggingface.co/meta-llama/CodeLlama-13b-Instruct-hf.
+Dataset. As mentioned in Section 3, four-fifth of the instances from SoAyBench are utilized here to train the SoAyLLaMA.
+Training Details.
+We train 13B models on 8 NVIDIA A100 GPUs (80GB) for 13 hours to reach a convergent loss, using a batch size of 8.
+Additionally, we train 7B models on 4 NVIDIA A100 GPUs (80GB) for 7 hours with a batch size of 16.
+All the training details are saved in the training log in our Github Repo [^4] https://anonymous.4open.science/r/SoAy-C025/readme.md.
+Appendix D SoAyBench
+To assess AMiner API-using capabilities, the foundational AMiner APIs for LLMs to invoke and a test set should be provided.
+However, given the dynamic and specialization nature of academic data, with academic entities-related information rapidly changing, maintaining a test set with static answers proves challenging.
+To address this challenge, we clone AMiner’s APIs at a specific point in time to create a static service, from which we create a corresponding static test set including 792 meaningful academic questions (one-fifth of the whole SoAyBench) and the corresponding ground truth conducted from the cloned APIs.
+We extract a subset of the comprehensive AMiner database, which includes over 66,000 entities such as persons and publications, at a specific point in time, September 23, 2023.
+This subset serves as our experimental evaluation environment, and we clone an API Service based on it.
+Consequently, the results returned by these APIs, given the same parameters, will remain consistent over time.
+This dataset and the API service provide a foundation for our subsequent construction of the test set.
+Appendix E Implementation on Other Scenarios
+Although SoAy is developed for the AMiner API, it can be seen as a general method for enabling large language models (LLMs) to use domain-specific APIs.
+To verify this generality, and also to prove that similar coupling phenomena, like those observed with the AMiner API, also exist in other domain-specific APIs, we deploy SoAy in two other scenarios that provide information-seeking APIs, OpenLibrary and CrossRef.
+Due to GPU resource limitations, we did not train separate models for these two new scenarios.
+Instead, we implemented corresponding SoAyGPT models to validate the feasibility of setting up SoAy in a new API domain.
+Details can be found at https://github.com/RUCKBReasoning/SoAy/tree/master/scenarios.
+Figure 7.
+The prompt is designed during the inference of SoAyGPT.
+Specifically: $P_{s}$ signifies the prompt for solution generation, which instructs the LLMs to generate a solution from the user’s natural language query.
+$P_{p}$ represents the prompt for code generation.
+This prompt is used to instruct the LLMs to generate code from the user’s query and the corresponding solution.
+Note that the code generation prompt is tailored for each solution in the solution library $L_{S}$.
+If the given solution is found in $L_{S}$, its specific $P_{p}$ is used; otherwise, a general $P_{p}$ is employed.
+This general prompt is composed by sampling (Query, Solution, Code) triplets from the training data for each solution in $L_{S}$.
+$P_{a}$ denotes the prompt for answer generation, used to instruct the LLM to generate a natural language answer based on the query and the code execution result.
+In the prompt content: All black text, except the ellipses, represents the original prompt content.
+Orange highlighted text includes annotations added for better understanding in the context of creating a prompt.
+Purple highlighted text indicates the content awaiting LLM output.
+[…]
+The prompt is designed during the inference of SoAyGPT.